### v1

# Planar Turán numbers of three configurations

Abstract
The planar Tuán number of $H$, denoted by $ex_{\mathcal{P}}(n,H)$, is defined as the maximum number of edges in an $n$-vertex $H$-free planar graph. The exact value of $ex_{\mathcal{P}}(n,H)$ remains a mystery when $H$ is large (for example, $H$ is a long path or a long cycle), while tight bounds have been established for many small planar graphs such as cycles, paths, $\Theta$-graphs and other small graphs formed by a union of them. One representative graph among such union graphs is $K_{1}+L$ where $L$ is a linear forest without isolated vertices. Previous works solved the cases when $L$ is a path or a matching. In this work, we first investigate the planar Turán number of the graph $K_{1}+L$ when $L$ is the disjoint union of a $P_{2}$ and $P_{3}$. Equivalently, $K_{1}+L$ represents a specific configuration formed by combining a $C_{3}$ and a $\Theta_{4}$. We further consider the planar Turán numbers of the all graphs obtained by combining $C_{3}$ and $\Theta_{4}$. Among the six possible such configurations, three have been resolved in earlier works. For the remaining three configurations (including $K_{1}+(P_{2}\dot{\cup}P_{3})$), we derive tight bounds. Furthermore, we completely characterize all extremal graphs for the remaining two of these three cases. Additionally, for two other unsolved cases involving $K_{1}+L$, we establish improved bounds.
Keywords: Planar Turán number, small graphs, union of triangles, extremal graphs
AMS subject classification 2020: 05C35.
1 Introduction
A graph $G$ is called $G$-free if it does not contain $G$ as a subgraph. The Turán number of a graph $G$, denoted by $ex(n,G)$, is the maximum number of edges in an $n$-vertex $G$-free graph. Turán-type problems are central topics in extremal combinatorics, employing diverse methodologies and intersecting with multiple mathematical disciplines. In 2016, Dowden [1] initiated the study of pl anar Turán-type problems. The planar Turán number of $H$, denoted by $ex_{\mathcal{P}}(n,H)$, is the maximum number of edges in an $n$-vertex $H$-free planar graph. Since planar graphs constitute a special and highly sparse graph class, the study of planar Turán problems relies primarily on structural methods.
Many planar graphs, particularly large and dense ones, have a planar Turán number of 3n63n-6. This trivial value is achieved because triangulations can be constructed to avoid specific forbidden subgraphs. For example, if $H$ contains at least three vertex-disjoint cycles, then $ex_{\mathcal{P}}(n,H)=3n-6$, as demonstrated by the triangulation $K_{2}+P_{n-2}$. Lan, Shi, and Song [10] provided sufficient conditions for a graph to have this trivial planar Turán number, a complete characterization remains unknown. One such condition is the maximum degree exceeds six (in other words, $ex_{\mathcal{P}}(n,H)=3n-6$ is $\Delta(H)\geq 7$). This makes the study of planar Turán numbers particularly interesting for two classes of graphs: those that are highly sparse and those that are small in terms of order.
It is well-known that the existence of Hamiltonian cycles in planar graph is a mystery, which results in a particularly interesting but challenging problem for determining the exact value of planar Turán numbers of long paths and long cycles. Shi, Walsh, and Yu [14] proved that $ex_{\mathcal{P}}(n,C_{k})\leq 3n-6-4^{-1}k^{\log_{3}2}n$ for large $k$. Combined with a lower bound by Győri, Varga and Zhu [7], the planar Turán number for long cycles $C_{k}$ is $(3-\Theta(k^{\log_{2}3}))n$. Li [13] also proved that $ex_{\mathcal{P}}(n,2C_{k})=(3-\Theta(k^{\log_{2}3}))n$. For a large planar graph $H$, if its planar Turán is not the trivial value 3n63n-6, then it must contain a long path or cycle since $\Delta(H)\leq 6$. Therefore, determining the planar Turán numbers of long paths and long cycles is an elementary problem.
Given the aforementioned challenges, researchers have shifted their focus toward determining the exact planar Turán numbers of small planar graphs. So far, tight bounds are known only for a few small planar graphs, including short cycles of length up to eight [1, 5, 6], paths of order at most eleven [8], and small $\Theta$-graphs [10] (we use $\Theta_{k}$ to denote a set of graphs that are obtained from a cycle by adding an additional edge joining two non-consecutive vertices. It is clear that $\Theta_{k}$ is a single graph if $k=4,5$). Other studies have examined structures composed of combined cycles, such as unions of triangles sharing vertices or edges [3, 11] and disjoint union of graphs [2, 11]. Among these configurations, one of the most interesting is $K_{1}+L_{t}$, where $L_{t}$ is a linear forest of $t$ vertices (the graph $K_{1}+L_{t}$ is obtained by joining a new vertex to all the vertices of $L_{t}$). Lan, Shi, and Song [9] proved that
Theorem 1.1 ([9]).
For an integer $4\leq t\leq 6$, let $H$ be a graph on $t$ vertices consisting of disjoint paths. Then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{13(t-1)n}{4t-2}-\frac{12(t-1)}{2t-1}$ for all $n\geq t+1$.
It is worth noting that not all the upper bounds obtained are tight. They further determined a tight upper bound for $ex_{\mathcal{P}}(n,K_{1}+2P_{2})$ when $n\geq 5$, and provided an improved upper bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$. Subsequently, Fang, Wang, and Zhai [3] established the tight bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$ and the tight bound of $ex_{\mathcal{P}}(n,K_{1}+P_{t})$ for $3\leq k\leq 6$. In this paper, we extend this line of research by studying graphs of the form $K_{1}+H$. Specifically, we establish a tight bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H=P_{2}\dot{\cup}P_{3}$ is a disjoint union of $P_{2}$ and $P_{3}$.
Theorem 1.2.
$ex_{\mathcal{P}}(n,K_{1}+(P_{2}\dot{\cup}P_{3}))\leq\frac{13n}{5}-\frac{26}{5}$ for all $n\geq 72$, with equality if $n=2\pmod{5}$.
Note that $K_{1}+(P_{2}\dot{\cup}P_{3})$ corresponds to a specific configuration formed by combining $C_{3}$ and $\Theta_{4}$. We further investigate all possible unions of $C_{3}$ and $\Theta_{4}$. It is clear that there are six different combinations illustrated in Figure 1 ($H_{4}=K_{1}+(P_{2}\dot{\cup}P_{3})$).
(a) $H_{1}$
(b) $H_{2}$
(c) $H_{3}$
(d) $H_{4}$
(e) $H_{5}$
(f) $H_{6}$
Figure 1: Six types of combinations of $C_{3}$ and $\Theta_{4}$.
Dowden [1] determined that $ex_{\mathcal{P}}(n,H_{1})=3n-6$ when $n\geq 6$. Fang, Wang, and Zhai [3] showed that $ex_{\mathcal{P}}(n,H_{2})\leq\frac{8}{3}(n-2)$ with a sharp upper bound. Because $H_{1}$ is a subgraph of $H_{3}$, $ex_{\mathcal{P}}(n,H_{3})=3n-6$ follows directly. In addition to $H_{4}$, we also focus on determining the planar Turán numbers of the remaining two graphs $H_{5}$, and $H_{6}$.
For $H_{5}$, we determine its planar Turán number exactly when $n=10x+6y$ has integer solutions $x\geq 2$ and $y\geq 0$, and characterize all extremal graphs.
Theorem 1.3.
$ex_{\mathcal{P}}(n,H_{5})\leq\lfloor\frac{5n}{2}\rfloor-4$ for all $n\geq 6$, with equality if $n=10x+6y$ has integer solutions $x\geq 2$ and $y\geq 0$. Moreover, the extremal graphs can be characterized (see Remark 1).
We determine the planar Turán number of $H_{6}$, which is the disjoint union of $C_{3}$ and $\Theta_{4}$, and characterize all extremal graphs. For convenience, we use $M_{t}$ to denote the union of $\left\lfloor\frac{t}{2}\right\rfloor$ pairwise vertex-disjoint edges and $\left\lceil\frac{t}{2}\right\rceil-\left\lfloor\frac{t}{2}\right\rfloor$ isolated vertices. To describe the extremal $H_{6}$-free planar graphs, we need the outerplanar Turán number of $C_{3}$ (denoted $ex_{\mathcal{OP}}(n,C_{3})$), which is the maximum number of edges in an $n$-vertex $C_{3}$-free outerplanar graph. Fang and Zhai [4] proved that $ex_{\mathcal{OP}}(n,C_{3})=\frac{3n-4}{2}$ for each even integer $n\geq 4$.
Theorem 1.4.
If $n\geq 174$, then $ex_{\mathcal{P}}(n,C_{3}\dot{\cup}\Theta_{4})=\left\lfloor\frac{5n}{2}\right\rfloor-4$. Moreover, if $n$ is even, then the extremal planar graph is a copy of $M_{n-2}+K_{2}$; if $n$ is odd, then the extremal planar graph is either a copy of $M_{n-2}+K_{2}$ or $O+\{u\}$, where $O$ represents a $C_{3}$-free outerplanar graph of even order with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
Let $L^{\prime}_{k}$ denote a linear forest of order $k$ without isolated vertices. Authors in [1, 3, 9] determined $ex_{\mathcal{P}}(n,K_{1}+L^{\prime}_{k})$ for all $L^{\prime}_{k}$ except for the two cases $L^{\prime}_{6}=P_{2}\dot{\cup}P_{4}$ and $L^{\prime}_{6}=2P_{3}$. We give a lower bound and an upper bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H\in\{P_{2}\cup P_{4},2P_{3}\}$, which offers a slight improvement to the upper bounds provided by Lan, Shi and Song (Theorem 1.1) for these scenarios.
Theorem 1.5.
Let $H\in\{P_{2}\cup P_{4},2P_{3}\}$, then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{41n}{14}-\frac{36}{7}$. In addition, if $12|n$, then $ex_{\mathcal{P}}(n,K_{1}+H)\geq\frac{11n}{4}-3$.
The rest of this paper is organized as follows. In Section 2, we introduce some concepts and notation on planar graphs. In Sections 3, 4 and 5, we establish sharp bounds on the planar Turán number of $H_{4},H_{5}$ and $H_{6}$, respectively, and characterize all extremal planar graphs for the latter two graphs. In Section 6, we give a short proof of Theorem 1.5.
2 Preliminaries
We first present essential definitions and preliminary results. A graph is planar if it can be drawn in the plane without edges crossing except at vertices. Such a drawing is a planar embedding, and a planar graph with a planar embedding is called a plane graph. A chord of a cycle $C$ in a graph $G$ is an edge not in $C$ whose endpoints both lie on $C$. For a face $F$ of a connected plane graph $G$, we use $\partial(F)$ to denote the boundary of $F$, which is a closed walk. If the length of this closed walk if $k$, then we call $F$ a $k$-face. Specifically, if $G$ is $2$-connected, then $\partial(F)$ is a cycle, and we call it the facial cycle of $F$. The outer boundary of $G$ is the boundary of its outer face. We always use $f_{k}(G)$ to denote the number of $k$-faces in the connected plane graph $G$.
A plane graph has a single unbounded outer face, with all other faces being inner faces. A near-triangulation is a $2$-connected plane graph where every inner face has degree three. A vertex with degree $k$ in $G$ is called a $k$-vertex. Let $E_{I}(G)$ be the set of all edges of $G$ that incident with two $3$-faces, and let $E_{I}(v)$ be the set of edges in $E_{I}(G)$ containing $v$ as an end vertex. For an edge $e=uv\in E_{I}(G)$, the union of its two adjacent 3-faces, say $F_{1}$ and $F_{2}$, forms a $\Theta$-graph of $uv$, denoted $\Theta_{uv}$ or $\Theta_{e}$. A $k$-wheel, denoted by $W_{k}$, is formed by connecting a single vertex to all vertices of a $k$-cycle. A $k$-fan is formed by connecting a single vertex to all vertices of a $k$-path. We refer to $K_{1}+tK_{2}$ as a friendship graph, where $t$ is an integer.
For a plane graph $G$ and two distinct inner 3-faces $F_{0}$ and $F_{\ell}$, we say $F_{0}$ is triangular-connected to $F_{\ell}$ (denoted $F_{0}\sim F_{\ell}$) if there exists an alternating sequence $F_{0}e_{1}F_{1}e_{2}\ldots e_{\ell}F_{\ell}$ such that $e_{i}$ is incident with both $F_{i}$ and $F_{i-1}$. For a 3-face $F$, let $\hat{F}$ be the set of inner 3-faces of $G$ that are triangular-connected to $F$. The triangular-block (TB for brief) $\hat{F}$] is the plane subgraph induced by the vertices and edges of the 3-faces in $\hat{F}$. Observe that TBs are edge-disjoint. Two TBs are adjacent if they share vertices in $G$, and we call such vertices junction vertices. A triangular-component (or shortly TC) is recursively constructed as follows:
  1. 1.
    Initialize with a TB $H=H_{0}$ of $G$.
  2. 2.
    Iteratively append any TB $H_{i}$ adjacent to $H$, updating $H=H\cup H_{i}$.
  3. 3.
    Terminate when no further adjacent TBs exist.
An inner face $F$ of the TB $B$ (resp. the TC $C$) that is not a face of $G$ is called a hole of $B$ (resp. a hole of $C$). Note that a hole in $B$ or $C$ may be a 3-face of $B$ or $C$ but not a 3-face of $G$. Let $S$ be a subgraph of $G$. We denote by $\Delta_{S}$ the number of inner 3-faces of $S$ that are also 3-faces of $G$. The triangle-density of $S$ is then defined as $\rho(S)=\frac{\Delta_{S}}{|S|}$. It is important to emphasize that only 3-faces of $S$ that are also 3-faces in $G$ contribute to the computation of $\Delta_{S}$.
For each $H$-free TB, we can transform it into a new TB by regarding all holes whose facial cycles are $C_{3}$s with corresponding 3-faces, such that the new TB maintains $H$-free. We call such a new TB a solid TB (i.e., a solid TB contains no holes with facial cycle $C_{3}$). In the proofs of Theorems 1.2 and 1.3, we will estimate the maximum triangle-density among all $H$-free TBs and TCs (where $H$ is either $H_{4}$ or $H_{5}$). For this purpose, it suffices to estimate the maximum triangle-density among all $H$-free solid TBs and all $H$-free TCs composed of solid TBs. See Figure 2 as an example, the left side is a plane graph, whose TBs are $B_{1},B_{2}$ and $B_{3}$ (see the Figure 2 (2), two gray faces are holes). $B_{1}^{\prime}$ is a solid TB obtained from $B_{1}$ (see the Figure 2 (3)).
Figure 2: An example on TBs and solid TBs.
3 Proof of Theorem 1.2
This section focuses on the planar Turán number of $H_{4}=K_{1}+(P_{2}\dot{\cup}P_{4})$. We begin with a proposition on $H_{4}$-free solid TBs, followed by a characterization of all such solid TBs.
Proposition 3.1.
Let $B$ be an $H_{4}$-free solid TB with $|B|\geq 4$ and outer boundary $C$. Then there exists a vertex $v\in V(C)$ such that BvB-v is a solid TB of order |B|1|B|-1, unless $B\cong B_{15}^{(n)}$ for even $n\geq 6$ (see Figure 3).
Proof.
The proof proceeds by considering two separate cases.
Case 1. $B$ is a near-triangulation.
If $C$ is chordless, then BvB-v is a solid TB of order |B|1|B|-1 for each $v\in V(C)$. Otherwise, each chord partitions $B$ into two solid TBs. Choose a chord $ab$ of $C$ such that the smaller solid TB, say $B^{\prime}$, has minimum order. Assume that the outer boundary of $B^{\prime}$ is $C^{\prime}$. Choose $v\in V(C^{\prime})-\{a,b\}$. By the chord selection criterion, $v$ is incident to no chord of $C$. Hence, BvB-v is a solid TB.
Case 2. $B$ is not a near-triangulation.
Let $\mathcal{F}$ be the set of holes in $B$. Since $B$ is a solid TB, for each $F\in\mathcal{F}$, $|V(\partial(F))|\geq 4$ and $|V(\partial(F))\cap V(C)|\leq 1$.
Case 2.1. For each face $F\in\mathcal{F}$, $|V(\partial(F))\cap V(C)|=0$.
We first introduce a useful fact that will be frequently utilized in the following proof. Note that the following result holds since all faces incident to $v\in V(C)$ are 3-faces.
Fact 1: For each $v\in V(C)$, $B[N(v)]$ is a path $P_{k}$, where $k=|N(v)|$.
Since $B$ is $H_{4}$-free, Fact 1 implies $2\leq|N(v)|\leq 4$ for each $v\in C$. If some $v\in V(C)$ has $d_{B}(v)=2$, say $N(v)=\{u,w\}$, then uvwuuvwu is a 3-face by Fact 1. Then $B-\{v\}$ is a solid TB, so we can assume $3\leq d_{B}(v)\leq 4$ for all $v\in V(C)$.
If $d_{B}(v)=3$ for all $v\in V(C)$, let $C=v_{1}v_{2}v_{3}\cdots v_{|C|}v_{1}$, and let $u$ be the third neighbor of $v_{1}$ (distinct from $v_{2}$ and $v_{|C|}$). Because $B[N[v_{1}]]$ is a 3-fan, $v_{|C|}uv_{2}$ is a path, so $u$ is adjacent to $v_{2}$. Inductively, since $B[N[v_{i-1}]]$ is a 3-fan for each $i\geq 3$, $u$ is adjacent to $v_{i}$. Thus, $B$ is the wheel graph $W_{|C|}$, implying $B$ is a near-triangulation, a contradiction. Therefore, $C$ must contain a vertex of degree 4.
We claim that $d_{B}(v)=4$ for each $v\in V(C)$. Suppose, for contradiction, that there exist adjacent vertices $u,v\in V(C)$ with $d_{B}(u)=3$ and $d_{B}(v)=4$. By Fact 1, $B[N(v)]$ is a path. Let $u,x_{1},x_{2},x_{3}$ denote the neighbors of $v$ listed in counter-clockwise order around $v$, and let $y_{1},x_{1},v$ be the neighbors of $u$ listed in counter-clockwise order around $u$. If $y_{1}=x_{3}$, then $x_{1}$ and $x_{3}$ are both neighbors of $u$, and $x_{1}x_{3}\in E(B)$. Since $d_{B}(x_{3})\leq 4$, it follows that $x_{1}x_{2}x_{3}x_{1}$ is a $3$-face of $B$. Hence, $B$ is a near-triangulation, a contradiction. So we assume $y_{1}\neq x_{3}$. Since no inner non-3-face intersects $C$ at any vertex, the edge $ux_{1}$ lies on two 3-faces: $vux_{1}v$ and $uy_{1}x_{1}u$. Observe that $y_{1}\in V(C)$ and $d_{B}(y_{1})\geq 3$, so $y_{1}x_{1}\notin E(C)$. Similarly, $y_{1}x_{1}$ lies on two 3-faces: $uy_{1}x_{1}u$ and $y_{1}y_{2}x_{1}y_{1}$, where $y_{2}\in V(B)$. If $y_{2}\neq x_{2}$, then $y_{1}y_{2}x_{1}y_{1}\cup\Theta_{vx_{1}}$ is an $H_{4}$, a contradiction. Thus, $y_{2}=x_{2}$. Now, consider $x_{3}\in V(C)$, and let $v,v^{\prime}$ be its neighbors in $C$. If $y_{1}=v^{\prime}$, then $B$ becomes a near-triangulation, a contradiction. Otherwise, $y_{1}\neq v^{\prime}$, and $x_{2}x_{3}$ is belongs to two $3$-faces of $B$. Consequently, $y_{1}x_{1}x_{2}y_{1}\cup\Theta_{x_{2}x_{3}}$ forms a $H_{4}$, yielding the final contradiction. Hence, every vertex in $V(C)$ has degree $4$ in $B$.
Let $v_{|C|},x_{1},x_{2},v_{2}$ be the neighbors of $v_{1}$ in counter-clockwise order. By Fact 1, $v_{|C|}x_{1}x_{2}v_{2}$ is a path. Similarly, for each $v_{i}\in C$, let $v_{i-1},x_{i-1},x_{i},v_{i+1}$ be its neighbors in counter-clockwise order (indices modulo $|C|$). By Fact 1, $v_{i-1}x_{i}x_{i+1}v_{i+1}$ is a path, so $x_{i}$ has neighbors $x_{i-1},v_{i},v_{i-1},x_{i+1}$. Now, suppose a vertex lies inside the cycle $C^{\prime}=x_{0}x_{1}\cdots x_{|C|-1}x_{0}$. Since $B$ is connected, some $x_{i}$ must have degree at least 5. Let $y$ be its fifth neighbor. The edge $x_{i}y$ lies in a 3-face of $B$, forcing $B[N[x_{i}]]$ to contain $H_{4}$, a contradiction. Thus, $V(B)=V(C)\cup V(C^{\prime})$ and $E(B)=E(C)\cup E(C^{\prime})\cup\{v_{i}x_{i},v_{i}x_{i+1}:i\in[|C|]\}$. In fact, the solid TB $B$ is isomorphic to $B_{15}^{(n)}$ for even $n\geq 6$, as shown in Figure 3.
Case 2.2. There exists a face $F\in\mathcal{F}$ such that $|V(\partial(F))\cap V(C)|=1$.
Assume that $V(\partial(F))\cap V(C)=\{v\}$. Suppose $v$ is incident to $a$ faces $F_{1},\ldots,F_{a}$ (in counter-clockwise order) that are not 3-faces. Since each edge of $B$ lies in at least one 3-face, there is a 3-face between any two adjacent faces $F_{i}$ and $F_{i+1}$. Because $B$ is $H_{4}$-free, $B[N[v]]$ is a friendship graph. Now we show that BvB-v is a solid TB. Let $F^{\prime}$ and $F^{\prime\prime}$ be two 3-faces in BvB-v, and let $\mathcal{P}$ be the set of connecting sequences between them in $B$. If some alternating sequence $P\in\mathcal{P}$ avoids 3-faces containing $v$, then $F^{\prime}\sim F^{\prime\prime}$ in BvB-v. Otherwise, every $P\in\mathcal{P}$ includes a 3-face $F$ containing $v$, and thus contains a sub-alternating sequence $F^{*}eFeF^{*}$, where $F^{*}$ is a 3-face with $F^{*}\cap F=\{e\}$. Removing all such sub-alternating sequences from $P$ yields a residual alternating sequence $P^{\prime}$ connecting $F^{\prime}$ and $F^{\prime\prime}$ within BvB-v. Therefore, $F^{\prime}$ and $F^{\prime\prime}$ are triangular-connected in BvB-v, implying BvB-v is a solid TB. ∎
Subsequently, we shall characterize all $H_{4}$-free solid TBs. This will be accomplished by initially analyzing the minimal configurations and then, in accordance with Proposition 3.1, systematically extending them through the addition of vertices. These $H_{4}$-free solid TBs are exhibited in Figure 3.
(a) $B_{1}$
(b) $B_{2}$
(c) $B_{3}$
(d) $B_{4}$
(e) $B_{5}$
(f) $B_{6}$
(g) $B_{7}$
(h) $B_{8}$
(i) $B_{9}$
(j) $B_{10}$
(k) $B_{11}^{(n)}$($n$ is even and $n\geq 4$)
(l) $B_{12}^{(n)}$ ($n$ is odd and $n\geq 5$)
(m) $B_{13}^{(n)}$($n$ is odd and $n\geq 7$)
(n) $B_{14}^{(n)}$ ($n$ is even and $n\geq 6$)
(o) $B_{15}^{(n)}$ ($n$ is even and $n\geq 6$)
Figure 3: $H_{4}$-free triangular blocks, with dashed circles indicating potential junction vertices. $|B_{i}^{(n)}|=n$ for $i\in\{11,12,13,14,15\}$.
Lemma 3.2.
Every $H_{4}$-free solid TB is isomorphic to a configuration in Figure 3.
Proof.
All solid TBs of order at most 5 are $H_{4}$-free, as exemplified by configurations $B_{1}$ to $B_{5}$, $B_{11}^{(4)}$, and $B_{12}^{(5)}$. Additionally, $B_{15}^{(n)}$ (see Figure 3) is $H_{4}$-free for even $n\geq 6$. Hence, we assume $B$ is $H_{4}$-free solid TB of order $n\geq 6$ and $B$ is not isomorphic to $B_{15}^{(n)}$ for any even $n\geq 6$.
For $|B|=6$, since $B$ is not isomorphic to $B_{15}^{(6)}$, Proposition 3.1 ensures that there is a vertex $v$ on the outer boundary of $B$ such that BvB-v is a copy of $B_{3}$, $B_{4}$, $B_{5}$ or $B_{12}^{(5)}$ as a subgraph. If BvB-v is a copy of $B_{3}$ or $B_{5}$, then $B$ contains an $H_{4}$ as a subgraph, which is impossible. If BvB-v is a copy of $B_{4}$, then $B$ is isomorphic to $B_{7}$, $B_{8}=B_{14}^{(6)}$ or $B_{9}$. If BvB-v is a copy of $B_{12}^{(5)}$, then $B$ is isomorphic to $B_{6}$, $B_{7}$, $B_{8}$ or $B_{11}^{(6)}$.
For $|B|=7$, Proposition 3.1 ensures that BvB-v is a copy of $B_{6}$, $B_{7}$, $B_{8}$, $B_{9}$, $B_{11}^{(6)}$ or $B_{15}^{(6)}$ (note $B_{9}\cong B^{6}_{15}$). If BvB-v is a copy of $B_{6},B_{8}$, $B_{9}$ or $B_{15}^{(6)}$, then $B$ contains an $H_{4}$, a contradiction. If BvB-v is a copy of $B_{7}$, then $B$ is isomorphic to $B_{10}$. If BvB-v is a copy of $B_{11}^{(6)}$, then $B$ is isomorphic to $B_{12}^{(7)}$ or $B_{13}^{(7)}$.
For $|B|=8$, since $B$ is not isomorphic to $B_{15}^{(8)}$, there is a vertex $v$ on the outer boundary of $B$ such that BvB-v is a copy of $B_{10}$, $B_{12}^{(7)}$, or $B_{13}^{(7)}$. If BvB-v is a copy of $B_{10}$ or $B_{13}^{(7)}$, then $B$ contains an $H_{4}$, a contradiction. If BvB-v is a copy of $B_{12}^{(7)}$, then $B$ is isomorphic to $B_{11}^{(8)}$ or $B_{14}^{(8)}$.
Now we assume that $|B|=n\geq 9$, since $B$ is not isomorphic to $B_{15}^{(n)}$ for any even $n\geq 10$, there is a vertex $v$ on the outer boundary of $B$ such that BvB-v is a solid TB of order n1n-1. If $n$ is odd, then BvB-v is a copy of $B_{11}^{(n-1)}$, $B_{14}^{(n-1)}$ or $B_{15}^{(n-1)}$. If BvB-v is a copy of $B_{11}^{(n-1)}$, then $B$ is isomorphic to $B_{12}^{(n)}$ or $B_{13}^{(n)}$. If $B$ is a copy of $B_{14}^{(n-1)}$ or $B_{15}^{(n-1)}$, then $B$ contains an $H_{4}$, a contradiction. If $n$ is even, then BvB-v is a copy of $B_{12}^{(n-1)}$ or $B_{13}^{(n-1)}$. If BvB-v is a copy of $B_{12}^{(n-1)}$, then $B$ is isomorphic to $B_{11}^{(n)}$, $B_{14}^{(n)}$. If BvB-v is a copy of $B_{13}^{(n-1)}$, then $B$ contains an $H_{4}$, a contradiction. ∎
Case $B_{1}$ $B_{2}$ $B_{3}$ $B_{4}$ $B_{5}$ $B_{6}$ $B_{7}$ $B_{8}$
$\Delta_{B_{i}}$ 1 3 4 4 5 4 5 6
Triangle density $1/3$ $3/4$ $4/5$ $4/5$ 1 $2/3$ $5/6$ 1
Case $B_{9}$ $B_{10}$ $B_{11}^{(n)}$ $B_{12}^{(n)}$ $B_{13}^{(n)}$ $B_{14}^{(n)}$ $B_{15}^{(n)}$
$\Delta_{B_{i}}$ 7 6 n2n-2 n2n-2 n1n-1 n1n-1 $n$
Triangle density $7/6$ $6/7$ (n2)/n(n-2)/n (n2)/n(n-2)/n (n1)/n(n-1)/n (n1)/n(n-1)/n 11
Table 1: The triangle-densities of triangle-blocks in $\mathcal{B}$.
Let $\mathcal{B}=\{B_{i}|1\leq i\leq 10\}\cup\{B^{(n)}_{i}|11\leq i\leq 15\}$. Table 1 summarizes the triangle-densities for solid TBs in $\mathcal{B}$. Let $D$ be an $H_{4}$-free TC composed of solid TBs. One can easily verify that $D$ is formed by connecting solid TBs in $\mathcal{B}$ via junction vertices (see Figure 3, where the junction vertices in each solid TB are marked by dots enclosed with dashed circles). Clearly, suppose, for contradiction, that there are two solid TBs $B_{1}$ and $B_{2}$ such that a non-junction vertex $v$ of $B_{1}$ is also a vertex in $B_{2}$. Let $F$ be an inner $3$-face of $B_{2}$ containing $v$. If $V(B_{1})\cap V(\partial(F))=\{v\}$, then $B_{1}\cup B_{2}$ contains an $H_{4}$ as a subgraph. For the case where $|V(B_{1})\cap V(\partial(F))|\geq 2$, we can similarly confirm that $B_{1}\cup B_{2}$ contains an $H_{4}$ through different combinations. Hence, in both cases, $B_{1}\cup B_{2}$ contains an $H_{4}$, a contradiction.
It is worth noting that not all solid TBs in $\mathcal{B}$ admit such a junction vertex; for example, the solid TBs in
$\mathcal{B}_{1}=\{B_{2},B_{4},B_{5},B_{8},B_{9},B_{15}^{(n)}\}$
are excluded due to the $H_{4}$-free constraint of $D$. For the same reason, some solid TBs contain at most one junction vertex, such as solid TBs in
$\mathcal{B}_{2}=\{B_{3},B_{7},B_{13}^{(n)},B_{14}^{(n)}\}.$
The remaining solid TBs are denoted by
$\mathcal{B}_{3}=\{B_{1},B_{6},B_{10},B_{11}^{(n)},B_{12}^{(n)}\},$
each containing at least two junction vertices. Define a solid TB $B$ as $B_{i}$-type TB if $B\cong B_{i}$. Before proceeding the proof of Theorem 1.2, we establish a foundational lemma on triangle-densities of $D$.
Lemma 3.3.
Let $D$ be an $H_{4}$-free TC of order at least seven. If $D\notin\{B_{15}^{(8)},B_{15}^{(10)}\}$, then $\rho(D)\leq\frac{6|D|-12}{5|D|}$.
Proof.
If we replace each TB in $D$ by the corresponding solid TB, then the resulting TC is also $H_{4}$-free, and its triangle density does not decrease. Hence, we can assume that $D$ is an $H_{4}$-free TC of order at least seven and is composed of solid TBs.
We proceed by induction on the number $k$ of solid TBs in $D$. For the base case $k=1$, since $|D|\geq 7$ and $D\notin\{B_{15}^{(8)},B_{15}^{(10)}\}$, $\rho(D)\leq\frac{6|D|-12}{5|D|}$ from Table 1. For the inductive step with $k\geq 2$, each solid TB must belong to $\mathcal{B}_{2}\cup\mathcal{B}_{3}$ as they are connected through junction vertices.
Case 1. There exists a solid TB of $D$, say $B$, containing exactly one junction vertex $v$.
Choose such a $B$ with $|B|$ minimum. It is easy to verify that $\Delta_{B}\leq|B|-1$ since $B\in\mathcal{B}_{2}\cup\mathcal{B}_{3}$. Let $D^{\prime}=D-V(B-v)$. If $|D^{\prime}|=3$, then both $D^{\prime}$ and $B$ are $B_{1}$-type TBs, and hence $\rho(D)=2/5<\frac{6|D|-12}{5|D|}$. Hence, we assume that $|D^{\prime}|\geq 4$. It is obvious that $D^{\prime}$ is a TC and $|D|=|D^{\prime}|+|B|-1$.
If $|D^{\prime}|\leq 6$, then $|D^{\prime}|\in\{4,5,6\}$. If $|D^{\prime}|=6$, then $\Delta_{D^{\prime}}\leq 5$, with equality holding only if $D^{\prime}\cong B_{7}$ or $D^{\prime}$ is the union of $B_{1}$ and $B_{6}$ by identifying three junction vertices. If $|D^{\prime}|=5$, then $\Delta_{D^{\prime}}\leq 4$, with equality holding only if $D^{\prime}\cong B_{3}$. If $|D^{\prime}|=4$, then $D^{\prime}\cong B_{11}^{4}$ and $\Delta_{D^{\prime}}=2$. In all these cases, $\Delta_{D^{\prime}}\leq|D^{\prime}|-1$. Therefore,
$\Delta_{D}\leq\Delta_{D^{\prime}}+(|B|-1)\leq(|D^{\prime}|-1)+(|B|-1)=|D^{\prime}|+|B|-2=|D|-1,$
and
$\rho(D)=\frac{\Delta_{D}}{|D|}\leq 1-\frac{1}{|D|}\leq\frac{6|D|-12}{5|D|},$
the last inequality holds since $|D|\geq 7$.
If $|D^{\prime}|\geq 7$, then by induction, $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. Note that $\Delta_{B}\leq|B|-1$. Hence,
$\Delta_{D}\leq\frac{6|D^{\prime}|-12}{5}+(|B|-1)<\frac{6(|D^{\prime}|+|B|-1)-12}{5}=\frac{6|D|-12}{5},$
implying $\rho(D)<\frac{6|D|-12}{5|D|}$.
Case 2. There is a solid TB of $D$, denoted $H$, containing precisely two junction vertices $u$ and $v$.
We can assume that each solid TB in $D$ belongs to $\mathcal{B}_{3}$, for otherwise there is a solid TB with only one junction $v$, which has been discussed in Case 1. Let $D^{\prime}$ be obtained form $D$ by removing all edges in $H$, and then deleting $V(H)-\{u,v\}$. Obviously, $|D^{\prime}|=|D|-|H|+2$.
Case 2.1. $D^{\prime}$ is connected.
If $|D^{\prime}|\geq 7$, then by induction, $\frac{\Delta_{D^{\prime}}}{|D^{\prime}|}\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. If $|D^{\prime}|\leq 6$, then either $D^{\prime}$ is a solid TB in $\{B_{1},B_{6},B_{11}^{(4)},B_{11}^{(6)},B_{12}^{(5)}\}$ or $D^{\prime}$ is one of the following configurations (see Figure4):
  1. 1.
    the union of two or three $B_{1}$-type TBs (see graphs (a) and (b) in Figure 4),
  2. 2.
    the union of $B^{4}_{11}$ and $B_{1}$ (see the graph (c) and (e) in Figure 4),
  3. 3.
    the union of two $B^{4}_{11}$-type TBs (see the graph (f) in Figure 4),
  4. 4.
    the union of $B^{5}_{12}$ and $B_{1}$ (see the graph (d) in Figure 4),
In all cases, one can verify that $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$.
Figure 4: all possible configurations of $D^{\prime}$ when $|D^{\prime}|\leq 6$, where the gray area denotes the hole.
If $H\ncong B_{10}$, then $\Delta_{H}\leq|H|-2$. Thus,
$\rho(D)=\frac{\Delta_{D^{\prime}}+\Delta_{H}}{|D|}\leq\frac{\frac{6|D^{\prime}|-12}{5}+|H|-2}{|D|}=\frac{6|D|-12-(|H|-2)}{5|D|}<\frac{6|D|-12}{5|D|}.$
If $H\cong B_{10}$, then
$\rho(D)=\frac{\Delta_{D^{\prime}}+6}{|D^{\prime}|+5}\leq\frac{\frac{6|D^{\prime}|-12}{5}+6}{|D|}=\frac{6(|D^{\prime}|+5)-12}{5|D|}=\frac{6(|D|)-12}{5|D|}.$
Case 2.2. $D^{\prime}$ is not connected.
Let $D^{\prime\prime}$ be obtained from $D$ by deleting $E(H)$ and then contracting $V(H)$ to a single vertex $w$. Then $D^{\prime\prime}$ is connected and $w$ is a cut-vertex of $D^{\prime\prime}$. Note that $w$ is a junction vertex of degree two in each solid TB. Hence, $D^{\prime\prime}$ remains a $H_{4}$-free TC. If $|D^{\prime\prime}|\geq 7$, then by induction, $\frac{\Delta_{D^{\prime\prime}}}{|D^{\prime\prime}|}\leq\frac{6|D^{\prime\prime}|-12}{5|D^{\prime\prime}|}$. If $|D^{\prime\prime}|\leq 6$, then since $|D^{\prime}|=|D^{\prime\prime}|+1$ and $D^{\prime}$ contains at least two components, it follows that $D^{\prime\prime}$ is a configuration of (c) or (e) in Figure 4, which implies that $\frac{\Delta_{D^{\prime\prime}}}{|D^{\prime\prime}|}\leq\frac{6|D^{\prime\prime}|-12}{5|D^{\prime\prime}|}$. Therefore, $\frac{\Delta_{D^{\prime\prime}}}{|D^{\prime\prime}|}\leq\frac{6|D^{\prime\prime}|-12}{5|D^{\prime\prime}|}$. Since $\Delta_{H}\leq|H|-1$ as $H\in\mathcal{B}_{3}$, it follows that
$\Delta_{D}\leq\frac{6|D^{\prime\prime}|-12}{5}+(|H|-1)=\frac{6|D|-12+(1-|H|)}{5}<\frac{6|D|-12}{5},$
implying $\rho(D)<\frac{6|D|-12}{5|D|}$.
Case 3. Each solid TB of $D$ possesses three junction vertices.
It is clear that each solid TB is either a copy of $B_{1}$ of a copy of $B_{6}$, and all three vertices of degree $2$ in the TB are junction vertices. Since each solid TB is $2$-connected, it follows that $D$ is $2$-connected. Hence, the boundary of each face in $D$ is a cycle. Thus, the inner faces of $D$ can be partitioned into two classes: the inner $3$-faces within each solid TB and holes. Note that if a hole is a $3$-face in $D$, then it is not a $3$-face in $G$.
Arbitrarily choose a hole $F$ (denoting its facial cycle by $C$), then each edge of $C$ is contained in exactly one solid TB of $D$. Furthermore, if a solid TB $B$ intersects $C$ in at least one edge, then:
  1. (A)
    if $B$ is isomorphic to $B_{1}$, then $|E(B)\cap E(C)|=1$, and
  2. (B)
    if $B$ is isomorphic to $B_{6}$, then $|E(B)\cap E(C)|=2$. Moreover, $B\cap C$ is a $3$-path whose endpoints are two of the junction vertices of $B$.
Case 3.1. $D$ has a hole $F$ whose facial cycle is $C=C_{3}$.
If each edge of $C$ belongs to a TB that is isomorphic to $B_{1}$ (say $V(C)=\{a_{1},a_{2},a_{3}\}$, and $a_{i}a_{i+1}$ is an edge of the $B_{1}$-type TB $T_{i}$), then by (A), $T_{1},T_{2}$ and $T_{3}$ are pairwise distinct. Since $D$ is $H_{4}$-free, the junction vertex $a_{i}$ is not in any other TBs. Let $D^{\prime}$ be a plane graph obtained from $D$ by deleting $a_{1},a_{2},a_{3}$. Then $D^{\prime}$ is an $H_{4}$-free TC obtained from $D$ by removing three solid TBs $T_{1},T_{2}$ and $T_{3}$. Additionally, since $D$ is $2$-connected and $a_{1},a_{2},a_{3}$ are not junction vertices, it follows that $D^{\prime}$ is connected. If $|D^{\prime}|\geq 7$, then by induction, $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. If $3\leq|D^{\prime}|\leq 6$, then $D^{\prime}$ is either a configuration of (a) or (b) in Figure 4, or a copy of $B_{6}$. In either cases, we can verify that $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. Therefore, $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$ holds. Since $\Delta_{D}-3=\Delta_{D^{\prime}}$ and $|D|-3=|D^{\prime}|$, it follows that
$\rho(D)\leq\frac{\Delta_{D^{\prime}}+3}{|D^{\prime}|+3}\leq\frac{\frac{6|D^{\prime}|-12}{5}+3}{|D^{\prime}|+3}=\frac{6(|D^{\prime}|+3)-15}{5(|D^{\prime}|+3)}<\frac{6|D|-12}{5|D|}.$
If there is an edge of $C$ belongs to a block $T$ that is isomorphic to $B_{6}$, then by (B), $|E(C)\cap E(T)|=2$ and $B\cap C$ is a $2$-path whose endpoints are two of the junction vertices of $T$. Without loss of generality, assume that $a_{1},a_{2},b$ are junction vertices of $T$ and $E(C)\cap E(T)=\{a_{1}a_{3},a_{2}a_{3}\}$. We further assume that $a_{1}a_{2}$ belongs to the block $T^{\prime}$. By (A), $T^{\prime}$ is ao $B_{1}$-type TB (assume that $V(T^{\prime})=\{a_{1},a_{2},c\}$). Let $D^{\prime}$ is a plane graph obtained from $D$ by deleting vertices $(V(T\cup T^{\prime})-\{b,c\})$ and removing $E(T)\cup E(T^{\prime})$. Since $D$ is $H_{4}$-free, the junction vertices $a_{1}$ and $a_{2}$ are not in any other TBs. Since $D$ is $2$-connected and $b,c$ are the only two junction vertices in $T\cup T^{\prime}$ connecting other solid TBs, it follows that $D^{\prime}$ is connected. If $|D^{\prime}|\geq 7$, then by induction, $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. If $3\leq|D^{\prime}|\leq 6$, then $D^{\prime}$ is either a configuration of (a) or (b) in Figure 4, or a copy of $B_{6}$. In either cases, we can verify that $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$. Therefore, $\rho(D^{\prime})\leq\frac{6|D^{\prime}|-12}{5|D^{\prime}|}$ holds. Since $\Delta_{D}-5=\Delta_{D^{\prime}}$ and $|D|-5=|D^{\prime}|$, it follows that
$\rho(D)\leq\frac{\Delta_{D^{\prime}}+5}{|D^{\prime}|+5}\leq\frac{\frac{6|D^{\prime}|-12}{5}+5}{|D^{\prime}|+5}<\frac{6|D|-12}{5|D|}.$
Case 3.2. Every hole of $D$ is not a $3$-face.
Note that we can add k3k-3 chords to each $k$-face of $D$ so that the final graph becomes a triangulation. Hence
$3n-6-e(D)\geq\sum_{F\in\mathcal{H}}(|F|-3),$
where $\mathcal{H}$ is the set of holes in $D$ and $|F|$ denote the number of vertices in the facial cycle of $F$. We denote by $m(D)=\sum_{F\in\mathcal{H}}(|F|-3)$ the number of all “missing edges” in $D$. To derive a lower bound of $m(D)$, we employ the discharging method. Initially, each hole $F$ in $D$ is assigned a charge of |F|3|F|-3, corresponding to its number of missing chords. This charge is then distributed equally among edges in the facial cycle of $F$. Suppose that $D$ has $t$ TBs isomorphic to $B_{1}$ and $s$ TBs isomorphic to $B_{6}$. Since all solid TBs of $B_{6}$ are connected only on junction vertices, it follows that $|D|\geq 3s+3$.
For a TB $T$ and each edge $e$ in the boundary of the outer face of $T$, let $F_{e}$ denote the unique hole in $D$ whose facial cycle contains $e$. Since each hole in $D$ is not a cycle, $|F_{e}|\geq 4$ and $e$ receives $\frac{|F_{e}|-3}{|F_{e}|}\geq\frac{1}{4}$ charge from $F_{e}$. Since a $B_{1}$-type TB has 3 boundary edges and a $B_{6}$-type TB has 6 boundary edges, the total charge received by boundary edges of $D$ is at least $\frac{1}{4}(3t+6s)$. It follows that $D$ can accommodate at least $3t/4+3s/2$ additional chords. Therefore,
$3|D|-6-e(D)\geq 3t/4+3s/2.$
While $e(D)=3t+9s$, so we obtain
$|D|\geq\frac{5t}{4}+\frac{7s}{2}+2.$
Thus,
$\rho(D)=\frac{\Delta_{D}}{|D|}\leq\frac{4s+t}{\frac{7s}{2}+\frac{5t}{4}+2}=\frac{16s+4t}{14s+5t+8}.$
When $s\geq 2$, we observe that
$\frac{16s+4t}{14s+5t+8}\leq\frac{8s}{7s+4}.$
Since $3s+3\leq|D|$ and $\frac{8s}{7s+4}$ is monotonically increasing with respect to $s$. We obtain
$\frac{8s}{7s+4}\leq\frac{\frac{8}{3}\left(|D|-3\right)}{\frac{7}{3}\left(|D|-3\right)+4}<\frac{6|D|-12}{5|D|}.$
When $s\leq 1$, since $|D|\geq 7$, we have that
$\rho(D)\leq\max\left\{\frac{4t}{5t+8},\frac{16+4t}{22+5t}\right\}<\frac{4}{5}\leq\frac{6|D|-12}{5|D|}.$
Now we are ready to prove the Theorem 1.2.
Proof of Theorem 1.2: Let $G$ be an $H_{4}$-free planar graph with $n\geq 72$ vertices that attains the maximum number of edges among all such graphs. Then $G$ must be connected. If $G$ is a triangulation, then there is a vertex $v$ of degree at least five, implying $G[N[v]]$ contains $H_{4}$ as a subgraph, a contradiction. Thus, $G$ admits a plane embedding where the outer boundary $\Gamma(G)$ is not a 3-face. In such an embedding, each 3-face is contained within a TB, which in turn is contained within a TC. Let $D_{1},D_{2},\cdots,D_{t}$ denote all the TCs of $G$, and let $\rho_{i}$ represent the triangle-density of $D_{i}$ for $i\in[t]$. If $D_{i}$ is isomorphic to $B_{1}$, $B_{6}$, or satisfies $|D_{i}|\geq 7$ while $D_{i}\notin\{B_{15}^{(6)},B_{15}^{(8)},B_{15}^{(10)}\}$, then by Lemma 3.3 and Table 1, we have:
$\rho_{i}\leq\frac{6|D_{i}|-12}{5|D_{i}|}.$
Moreover, the function $\frac{6|D_{i}|-12}{5|D_{i}|}$ is monotonically increasing in $|D_{i}|$. Consequently, we have $\rho_{i}\leq\frac{6n-12}{5n}$. For the remaining cases, $D_{i}$ is isomorphic to one of the following:
$\{B_{2},B_{3},B_{4},B_{5},B_{7},B_{8},B_{9},B^{(4)}_{11},B^{(6)}_{11},B^{(5)}_{12},B^{(6)}_{14},B^{(6)}_{15},B^{(8)}_{15},B^{(10)}_{15}\}$
(it is worth noting that $B^{(6)}_{14}=B_{8}$ and $B_{15}^{(6)}=B_{9}$). In these cases, a straightforward verification shows that $\rho_{i}\leq\frac{6n-12}{5n}$ holds for $n\geq 72$. Since $D_{i}$s are pairwise vertex-disjoint, $\sum_{i\in[t]}{|D_{i}|}\leq n$. Hence,
$f_{3}(G)=\sum_{i\in[t]}{|D_{i}|\cdot d_{i}}\leq\frac{6n-12}{5n}\sum_{i\in[t]}{|D_{i}|}\leq\frac{6n-12}{5}.$
Furthermore, we have
$2e(G)=\sum if_{i}(G)\geq 3f_{3}(G)+4(f(G)-f_{3}(G))=4f(G)-f_{3}(G).$
Combining with Euler’s formula e(G)f(G)+2=ne(G)-f(G)+2=n, we obtain
$e(G)\leq 2n-4+\frac{1}{2}f_{3}(G).$
Substituting the bound for $f_{3}(G)$ yields:
$e(G)\leq 2n-4+\frac{1}{2}f_{3}(G)=\frac{13n}{5}-\frac{26}{5}.$
To demonstrate the tightness of our inequality, consider the graph family $\{G_{k}:k\geq 14\}$ in Figure 5. It is clear that $G_{k}$ is a $H_{4}$-free plane graph with $|G_{k}|=4k+2$ and $e(G_{k})=\frac{13n}{5}-\frac{26}{5}$. This establishes the sharpness of our upper bound and completes the proof of Theorem 1.∎
$k$
Figure 5: The extremal graph $G_{k}$.
4 Proof of Theorem 1.3
This section is dedicated to studying the planar Turán number of $H_{5}$. We begin by introducing a structural property of $H_{5}$-free solid TBs, then fully characterize these solid TBs, and finally determine the triangle-densities of $H_{5}$-free TCs.
Proposition 4.1.
Let $B$ denote an $H_{5}$-free solid TB with $|B|\geq 4$, and let $C$ represent its outer boundary. Then, within the vertex set $V(C)$ or among the holes of $B$, there exists a vertex $v$ such that BvB-v is a solid TB of order |B|1|B|-1.
Proof.
As discussed in Proposition 3.1, the result holds when $B$ does not contain holes. Hence, assume that $B$ contains a hole. Let $\mathcal{F}$ be the set of holes of $B$. Then for each $F\in\mathcal{F}$, $|V(\partial(F))\cap V(C)|\leq 1$. Since $B$ is a solid TB, each $\partial(F)$ is a cycle of order at least four.
Case 1. For each $F\in\mathcal{F}$, $|V(\partial(F))\cap V(C)|=0$.
We employ Fact 1 in Proposition 3.1, which still hold in this case. If there exists a vertex $v\in V(C)$ with degree two, then BvB-v remains a solid TB. Thus, we assume that each vertex in $C$ has degree at least three. If there is a vertex $v\in V(C)$ such that $d(v)=t\geq 4$, then let $x_{1},x_{2},\ldots,x_{t}$ be neighbors of $v$ listed in counter-clockwise order around $v$ such that $x_{1},x_{t}\in V(C)$. Note that either $x_{1}x_{t-1}\notin E(B)$ or $x_{t}x_{2}\notin E(B)$. Hence, either $\Theta_{x_{1}x_{2}}\cup vx_{t-1}x_{t}v$ or $\Theta_{x_{t-1}x_{t}}\cup vx_{1}x_{2}v$ is an $H_{5}$, a contradiction. Thus, each vertex in $C$ is a $3$-degree vertex. Let $C=v_{1}v_{2}v_{3}\cdots v_{|C|}v_{1}$, and let $u$ denote the third neighbor of $v_{1}$ (distinct from $v_{2}$ and $v_{|C|}$). Since $v_{|C|}uv_{2}$ is a path, $u$ is adjacent to $v_{2}$. Proceeding inductively, for each $i\in[|C|]$, ‘we have that $u$ is adjacent to $v_{i}$. Consequently, $B$ is a wheel graph $W_{|C|}$, contradicting $B$ contains a hole.
Case 2. there exists a face $F\in\mathcal{F}$ such that $|V(\partial(F))\cap V(C)|=1$.
Let $v\in V(C)$ be a vertex incident to $a$ non–3-faces, denoted by $F_{1},\ldots,F_{a}$ in counter-clockwise order around $v$. Since each edge of $B$ lies in at least one 3-face, there is a 3-face between any two adjacent faces $F_{i}$ and $F_{i-1}$ for $i\in[a]$ (for convenience, we define $F_{0}$ as the outer face of $B$). We aim to show that either BvB-v is a solid TB or BuB-u is a solid TB for some $u\in\partial(F)$, where $F$ is a hole of $B$.
First, we claim that $B[N[v]]$ is a friendship graph. Otherwise, assume there exists an $i\in[a]$ such that there are $r\geq 2$ 33-faces between $F_{i-1}$ and $F_{i}$. Say $x_{1}x_{2}\ldots x_{r+1}$ are consecutive neighbors of $v$ listed in counter-clockwise order around $v$ such that $x_{1}\in V(\partial(F_{i-1}))$ and $x_{r+1}\in V(\partial(F_{i}))$, where $i$ take mudola $a$. Similarly, assume that $x^{\prime}_{1}x^{\prime}_{2}\ldots x^{\prime}_{\ell+1}$ be consecutive neighbors of $v$ listed in counter clockwise order around $v$ such that $x^{\prime}_{1}\in V(\partial(F_{i-2}))$ and $x^{\prime}_{\ell+1}\in V(\partial(F^{\prime}_{i-1}))$. Then $vx_{i}x_{i+1}v$ and $vx^{\prime}_{j}x^{\prime}_{j+1}v$ are $3$-faces of $B$ for $i\in[r]$ and $j\in[\ell]$. Let $y_{1}$ be the neighbor of $x_{1}$ on $F_{i-1}$ other than $v$. Note that $x_{1}y_{1}$ is in the unique $3$-face of $B$, say $x_{1}y_{1}z_{1}x_{1}$. We will complete the proof by several cases as follows.
  1. (1)
    $z_{1}=x_{2}$.
    Let $y_{2}$ be the neighbor of $y_{1}$ on $\partial(F_{i-1})$ such that $y_{2}\neq x_{1}$. Note that $y_{1}\neq x^{\prime}_{\ell+1}$ since $F_{i-1}$ is not a $3$-face. Without loss of generality, assume that $F^{*}$ is the unique $3$-face of $B$ containing $y_{1}y_{2}$, say $\partial(F^{*})=y^{\prime}y_{1}y_{2}y^{\prime}$. If $x^{\prime}_{\ell}=y_{1}$, then $y_{2}=x^{\prime}_{\ell+1}$ and $y^{\prime}=v$. One can verify that $B-y_{2}$ is a desired solid TB. If $x^{\prime}_{\ell}\neq y_{1}$, then $\Theta_{x_{1}x_{2}}\cup vx^{\prime}_{\ell}x^{\prime}_{\ell+1}v$ is a copy of $\Theta_{4}\cup C_{3}$, a contradiction.
  2. (2)
    $z_{1}=x_{3}$.
    Since $x^{\prime}_{\ell}\neq x_{i}$ and $x^{\prime}_{\ell+1}\neq x_{i}$ for each $i\in[r+1]$, it follows that $vx_{1}x_{2}v\cup vx^{\prime}_{\ell}x^{\prime}_{\ell+1}v\cup\{x_{2}x_{3},x_{1}x_{3}\}$ is a copy of $\Theta_{4}\cup C_{3}$, a contradiction.
  3. (3)
    $z_{1}\notin\{x_{2},x_{3}\}$.
    Then $x_{1}y_{1}z_{1}x_{1}\cup\Theta_{vx_{2}}$ is a copy of $\Theta_{4}\cup C_{3}$, a contradiction.
Therefore, $B[N[v]]$ is a friendship graph.
Now we show that BvB-v is a solid TB. This proof is similar to the proof in Case 2 of Proposition 3.1, and is restated here for convenience. Let $F^{\prime}$ and $F^{\prime\prime}$ be two 3-faces in BvB-v, and let $\mathcal{P}$ be the set of connecting sequences between them in $B$. If some alternating sequence $P\in\mathcal{P}$ avoids 3-faces containing $v$, then $F^{\prime}\sim F^{\prime\prime}$ in BvB-v. Otherwise, every $P\in\mathcal{P}$ includes a 3-face $F$ containing $v$, and thus contains a sub-alternating sequence $F^{*}eFeF^{*}$, where $F^{*}$ is a 3-face with $F^{*}\cap F=\{e\}$. Removing all such sub-alternating sequences from $P$ yields a residual alternating sequence $P^{\prime}$ connecting $F^{\prime}$ and $F^{\prime\prime}$ within BvB-v. Therefore, $F^{\prime}$ and $F^{\prime\prime}$ are triangular-connected in BvB-v, implying BvB-v is a solid TB. ∎
Lemma 4.2.
Let $B$ be an $H_{5}$-free solid TB. Then $B$ is isomorphic to one of the following configurations: (i) configurations $B_{1}$-$B_{6}$, $B_{11}^{(4)}$, and $B_{12}^{(5)}$ in Figure 3; (ii) $k$-wheel $W_{k}$ and $k$-fan $F_{k}$ ($k\geq 5$); (iii) the three exceptional configurations $B_{1}^{\prime},B_{2}^{\prime}$ and $B_{3}^{\prime}$ illustrated in Figure 6.
Proof.
All solid TBs of order at most 5 are $H_{5}$-free, as demonstrated by the structures $B_{1}$ to $B_{5}$, $B_{11}^{(4)}$, $B_{12}^{(5)}$ in Figure 3. If $|B|\geq 6$, Proposition 4.1 ensures there exists a vertex $v\in\partial(F)$ such that BvB-v is a solid TB, where $F$ is either the outer face of $B$ or a hole of $B$.
(a) $B^{\prime}_{1}$
(b) $B^{\prime}_{2}$
(c) $B^{\prime}_{3}$
Figure 6: A part of $H_{5}$-free triangle-blocks.
For $|B|=6$, $B-v\in\{B_{3},B_{4},B_{5},B_{12}^{(5)}\}$. If BvB-v is a copy of $B_{3}$, then $B\cong B^{\prime}_{1}$; if BvB-v is a copy of $B_{4}$, then $B$ contains $H_{5}$ as a subgraph, a contradiction; if BvB-v is a copy of $B_{5}$, then $B\cong B^{\prime}_{2}$; if BvB-v is a copy of $B_{12}^{(5)}$, then $B\in\{F_{5},W_{5},B_{6}\}$.
For $|B|=7$, BvB-v is a solid TB belonging to $\{B_{6},B^{\prime}_{1},B^{\prime}_{2},W_{5},F_{5}\}$. If BvB-v is a copy in $\{B_{6},B^{\prime}_{2},W_{5}\}$, then $B$ contains $H_{5}$ as a subgraph, a contradiction; If BvB-v is a copy of $B^{\prime}_{1}$, then $B\cong B^{\prime}_{3}$; If BvB-v is a copy of $F_{5}$, then $B\in\{F_{6},W_{6}\}$.
For $|B|=8$, BvB-v is a solid TB in $\{B^{\prime}_{3},W_{6},F_{6}\}$. If BvB-v is a copy of $B^{\prime}_{3}$ or $W_{6}$, then $B$ must contain $H_{5}$ as a subgraph, a contradiction. If BvB-v is a copy of $F_{6}$, then $B$ is isomorphic to one solid TB in $\{F_{7},W_{7}\}$.
For $|B|=9$, BvB-v is a solid TB with $B-v\in\{W_{7},F_{7}\}$. If BvB-v is a copy of $W_{7}$, then $B$ must contain $H_{5}$ as a subgraph, a contradiction. If BvB-v is a copy of $F_{7}$, then $B$ is isomorphic to one solid TB in $\{F_{8},W_{8}\}$;
Through inductively construction, we establish that for $|B|=k+1$ with $k\geq 9$, $B\in\{F_{k},W_{k}\}$. This completes this proof. ∎
For convenience, let $\mathcal{F}$ denote the set of all $H_{5}$-free solid TBs, then
$\mathcal{F}=\{B_{1},B_{2},B_{3},B_{4},B_{5},B_{6},B^{(4)}_{11},B^{(5)}_{12},B^{\prime}_{1},B^{\prime}_{2},B^{\prime}_{3}\}\cup\{W_{k},F_{k}:k\geq 5\}$.
Lemma 4.3.
If $D$ is an $H_{5}$-free TC consisting of solid TBs, then the triangle-density of $D$ is at most $1$. Moreover, the triangle-density of $D$ is one only if $D$ is a copy of $B_{5}$ or $B_{2}^{\prime}$.
Proof.
Table 2 shows triangle-densities of $B_{1}^{\prime},B_{2}^{\prime},B_{3}^{\prime},W_{n}$ and $F_{n}$. Combining Table 1, we have that the triangle-density of $D$ is at most $1$ when $D$ is a solid TB. Moreover, the triangle-density of $D$ is one only if $D\in\{B_{5},B_{2}^{\prime}\}$. Next, we assume that $D$ consists of at least two at least TBs. Our aim is to show that $\rho(D)<1$. For two solid TBs $B^{\prime},B^{\prime\prime}$ of $D$ with $|V(B^{\prime})\cap V(B^{\prime\prime})|\geq 1$, there is a $3$-face $F$ of $B^{\prime\prime}$ such that $|V(B^{\prime})\cap V(\partial(F))|\geq 1$. Choose such solid TBs $B^{\prime},B^{\prime\prime}$ and a $3$-face $F$ of $B^{\prime\prime}$ such that $|V(B^{\prime})\cap V(\partial(F))|$ is maximum.
Case 1. $|V(B^{\prime})\cap V(\partial(F))|=1$.
By the choice of $B^{\prime},B^{\prime\prime}$ and the face $F$ of $B^{\prime\prime}$, we have that $B^{\prime},B^{\prime\prime}$ are fan graphs, and $V(B^{\prime})\cap V(B^{\prime\prime})=\{v\}$ is the common center vertex of them. If there is a TB $B=F_{k}$ for some $k\geq 3$, then the center vertex of $B$, say $u$, is a cut-vertex of $D$. Let $D^{\prime}=D-(V(B)-v)$. By induction, $\rho(D^{\prime})\leq 1$. Since $\Delta_{B}=|B|-2$ and $|D|=|D^{\prime}|+|B|-1$, it follows that $\rho(D)<1$. Thus, we assume that each TB of $D$ is an $F_{2}=B_{1}$. Let $\mathcal{F}$ denote the set of all faces of $D$ not in any TBs. For each $F\in\mathcal{F}$, let $\ell_{F}$ denote the length of $\partial(F)$ (note that $\partial(F)$ is a closed trail). If there is a face $F$ such that $F$ is a $3$-face, then since $D$ is $H_{5}$-free, $D$ consists of three $B_{1}$s. Consequently, $\rho(D)=1/2<1$. Now, we assume that each $F\in\mathcal{F}$ is not a $3$-face. Note that
$e(D)+\sum_{F\in\mathcal{F}}(\ell_{F}-3)=3n-6.$ (1)
Assign a charge of $\ell_{F}-3$ to each $F\in\mathcal{F}$, and then distribute this charge equally among vertices in $\partial(F)$ (since $\partial(F)$ is a closed trail, the charge on some vertices may be counted more than once). Since each $F\in\mathcal{F}$ is not a $3$-face, if a vertex $v\in\partial(F)$ appears $k$ times in $\partial(F)$, then $v$ receives $k(\ell_{F}-3)/\ell_{F}\geq k/4$ charge. For each vertex of $v\in V(D)$, assume that there are $d_{v}$ TBs incident with $v$. Then $v$ receives $d_{v}(\ell_{F}-3)/\ell_{F}\geq d_{v}/4$ charge. Therefore,
$\sum_{v\in V(D)}d_{v}=e(D)$
and
$\sum_{F\in\mathcal{F}}(\ell_{F}-3)=\sum_{v\in V(D)}d_{v}(\ell_{F}-3)/\ell_{F}\leq\frac{1}{4}\sum_{v\in V(D)}d_{v}=\frac{1}{4}e(D).$
Combining with Ineq. (1), we have that $e(D)\leq(12|D|-24)/5$. Since each TB in $D$ is a $B_{1}$-type TB and any two TBs are edge-disjoint, it follows that there are at most $e(D)/3<|D|$ TBs in $D$. Hence, $\rho(D)=\Delta(D)/|D|<1$.
Case 2. $|V(B^{\prime})\cap V(\partial(F))|\geq 2$.
Then $B^{\prime}\in\{B_{11}^{(4)},B_{4},F_{4},F_{5},B_{6}\}$ and $B^{\prime}\cup\partial(F)$ is a graph as shown in Figure 7 (a)–(e). respectively; otherwise $B^{\prime}\cup B^{\prime\prime}$ contains an $H_{5}$, a contradiction. Since $D$ is $H_{5}$-free, we have that $B^{\prime\prime}$ is either a $B_{1}$-type TB (see Figure 7 (a)–(e)) or a $B_{11}^{(4)}$-type TB (see Figure 7 (A)–(E)), and $D=B^{\prime}\cup B^{\prime\prime}$. Consequently, the triangle-density of $D$ is less that $1$. ∎
Figure 7: All possible TBs of $B^{\prime}$.
Cases $B^{\prime}_{1}$ $B^{\prime}_{2}$ $B^{\prime}_{3}$ $W_{k}$ $F_{k+1}$
$\Delta_{B}$($B\in\{B^{\prime}_{i}(i\in[3]),W_{k},F_{k+1}$) $\leq$ 5 6 6 $k$ $k$
Triangle density $\leq$ $\frac{5}{6}$ 1 $\frac{6}{7}$ $\frac{k}{k+1}$ $\frac{k}{k+2}$
Table 2: The triangle-densities of a part of $H_{5}$-free triangle-components.
Now we are ready to give the proof of Theorem 1.3.
Proof of Theorem 1.3 Let $G$ be an $H_{5}$-free plane graph with $n\geq 6$ vertices and the maximum number of edges, which ensures $G$ is connected. If $G$ is a triangulation, then $G$ itself is an $H_{5}$-free TB, implying $G$ is a copy of $B_{1},B_{2}$ or $B_{5}$. This contradicts $n\geq 6$. Hence, we can assume that $G$ is embedded in the plane such that its outer face is not a 3-face. Let $D_{1},D_{2},\ldots,D_{t}$ denote all TCs in $G$, and let $\rho_{i}$ represent the triangle-density of $D_{i}$. Since the outer boundary of $G$ is not a $3$-face, each $3$-face is an inner face of some TB in $G$. Hence,
$f_{3}(G)=\sum_{i\in[t]}\Delta_{D_{i}}=\sum_{i\in[t]}|D_{i}|\rho(D_{i})\leq\sum_{i\in[t]}|D_{i}|\leq n,$ (2)
Then,
$2e(G)=\sum_{i\geq 3}if_{i}(G)\geq 3f_{3}(G)+4(f(G)-f_{3}(G))=4f(G)-f_{3}(G).$ (3)
Combining with Euler’s formula e(G)f(G)+2=ne(G)-f(G)+2=n, we obtain
$e(G)\leq 2n-4+\frac{1}{2}f_{3}(G).$
Thus,
$e(G)\leq 2n-4+\frac{1}{2}f_{3}(G)=\frac{5n}{2}-4.$ (4)
To demonstrate the sharpness of the inequality, let $k\geq 4$ be an even integer and $R^{\prime}_{k}$ be the plane graph shown in Figure 8, constructed from $k$ disjoint $B_{5}$ copies augmented with two cycles $C_{k}:=v_{1}v_{2}\cdots v_{k}v_{1}$ and $C_{2k}:=u_{1}u_{2}\cdots u_{2k}u_{1}$. Let $R_{k}$ be derived from $R^{\prime}_{k}$ by adding edge $\{v_{i}v_{k+1-i}:i\in[k/2]\}\cup\{u_{i}u_{2k+1-i}:i\in[k]\}$. It is clear that $R_{k}$ is an $H_{5}$-free plane graph.
Now we construct an extremal graph when $n=10x+6y$ has integer solutions $x\geq 2$ and $y\geq 0$. Let $H_{0}=R_{x}$ and let $H_{i}$ be a plane graph obtained from $H_{i-1}$ by adding a copy of $B_{2}^{\prime}$, say $B$, in a $4$-face $F$ of $H_{i-1}$, and then joining each vertex of $V(\partial(F))$ to a vertex on the outer boundary of $B$ such that the four new edges forms a matching. Then $H_{y}$ is an $n$-vertex $H_{5}$-free plane graph. ∎
Remark 1. Note that $e(G)=\frac{5n}{2}-4$ holds if and only if all equalities in (2), (3) and (4) hold, which implies
  1. (i)
    each $D_{i}$ is a copy of $B_{5}$ or $B_{2}^{\prime}$,
  2. (ii)
    $\bigcup_{i\in[t]}V(D_{i})=V(G)$, and
  3. (iii)
    each face of $G$ is either a $3$-face or a $4$-face.
Therefore, if $n=10x+6y$ has integer solutions $x\geq 2$ and $y\geq 0$, then each graph satisfies conditions (i), (ii) and (iii). The graph $H_{y}$ constructed above implies that such graphs exist definitely. ∎
$v_{1}$$u_{1}$$u_{2}$$v_{2}$$u_{3}$$u_{4}$$v_{k}$$u_{2k-1}$$u_{2k}$$C_{k}$$C_{2k}$
Figure 8: The graph $R^{\prime}_{k}$
5 Proof of Theorem 1.4
In this section, we study the planar Turán number of $C_{3}\dot{\cup}\Theta_{4}$. Let $G$ be a $C_{3}\dot{\cup}\Theta_{4}$-free plane graph with $|G|\geq 174$. A set of edges is called independent edges if they form a matching. The proof of Theorem 1.4 could be proceeded using the idea of analyzing the size of $E_{I}(G)$ proposed in [12]. The following result is necessary to avoid $C_{3}\dot{\cup}\Theta_{4}$ in $G$.
Lemma 5.1.
For any two independent edges $e$, $f$ in $E_{I}(G)$, $|V(\Theta_{e})\cap V(\Theta_{f})|\geq 2$.
Let $e=uv$ be an edge in $E_{I}(G)$ and $V(\Theta_{e})=\{u,v,x,y\}$. Define $A_{e}^{*}=\{\text{$e\in E_{I}(G)$: $e$}$ is incident to at least one vertex in $V(\Theta_{e})\}$ and let $A_{e}=E_{I}(G)\setminus A_{e}^{*}$. We provide the following observation.
Observation 5.2.
If $f\in A_{e}$, then $\Theta_{e}\cup\Theta_{f}$ must be isomorphic to one of the structures $\{D_{i}:i\in[3]\}$ illustrated in Figure 9.
eeffuuvvaabbxxyy
(a) $D_{1}$
eeffuuvvaabbxxyy
(b) $D_{2}$
eeffuuvvaabbxxyy
(c) $D_{3}$
Figure 9: The planar structures constituted by $\Theta_{e}\cup\Theta_{f}$ are $D_{1}$, $D_{2}$, and $D_{3}$. Specifically, $D_{1}$ contains two holes, namely auyvaauyva and xubvxxubvx. $D_{2}$ contains two holes, which are xvyaxxvyax and xuybxxuybx. And $D_{3}$ contains two holes, axuaaxua and buyvxbbuyvxb.
Let $\{e_{1},\cdots,e_{t}\}\subseteq E_{I}(G)$. For any inner face $F$ of the plane subgraph $H=\cup_{i\in[t]}\Theta_{e_{i}}$, if $F$ is not a 3-face in any $\Theta_{e_{i}}$, then $F$ is call a pseudo face of $H$. An edge or a vertex is said to lie in a pseudo face when it is contained in the interior region of the face’s closed boundary in the plane. Given a plane subgraph $P$ of $G$ and an edge subset $S=\{e_{1},\ldots,e_{t}\}\subseteq E_{I}(G)$, the generating graph of $P$ by $S$ is $P^{\prime}=P\cup\bigcup_{i=1}^{t}\Theta_{e_{i}}$. If $S$ consists of a single edge $e$, we say $P^{\prime}=P\cup\Theta_{e}$ is the generating graph of $P$ by $e$. Let $d_{I}(v)=d_{G[E_{I}(G)]}(v)$ and $\Delta_{I}(G)=\Delta(G[E_{I}(G)])$. We now establish a key lemma that will be essential for proving Theorem 1.4.
Lemma 5.3.
Let $G$ be a $C_{3}\dot{\cup}\Theta_{4}$-free planar graph with order $n\geq 174$. If $|E_{I}(G)|>\frac{n}{2}$ and $\Delta_{I}(G)\leq 9$, then $|E_{I}(G)|\leq\frac{n}{2}+4$. Moreover, if $|E_{I}(G)|=\left\lfloor\frac{n}{2}\right\rfloor+4$, then $G$ contains a subgraph that is isomorphic to $\left(\left\lfloor\frac{n-2}{2}\right\rfloor K_{2}\right)+K_{2}$.
Proof.
If $|E_{I}(G)|\leq 90$, then since $n\geq 174$, we derive $|E_{I}(G)|<\left\lfloor\frac{n}{2}\right\rfloor+4$, and the result follows immediately. We therefore proceed under the assumption that $|E_{I}(G)|>90$. Since $\Delta_{I}(G)\leq 9$, for each $e\in E_{I}(G)$ and each $f\in A_{e}$, $A_{e}$ must contain at least 4 independent edges including $f$. Based on the structure of $\Theta_{e}\cup\Theta_{f}$ shown in Observation 5.2, our analysis proceeds by examining three distinct cases.
Case 1. There exists an $e\in E_{I}(G)$ and an $f\in A_{e}$ such that $\Theta_{e}\cup\Theta_{f}$ is a copy of $D_{1}$.
Without loss of generality, let $f=ab$, $e=uv$ and $V(\Theta_{e})=\{u,v,x,y\}$. Then $V(\Theta_{f})=\{u,v,a,b\}$. Then the two pseudo faces of $\Theta_{e}\cup\Theta_{f}$ are $F_{1}=buxvb$ and $F_{2}=auyva$. Since $x,y$ belong to $\partial(F_{1})$ and $\partial(F_{2})$, respectively, it follows that $xy\notin E(G)$. Define $B_{e}^{*}$ as the set of edges in $E_{I}(G)$ that are incident with either $u$ or $v$, and let $B_{e}=E_{I}(G)-B^{*}_{e}$. Then $f\in B_{e}$.
Claim 5.4.
If $f^{\prime}=a^{\prime}b^{\prime}$ is an edge of $B_{e}$ such that $\{x,y\}\cap\{a^{\prime},b^{\prime}\}\neq\emptyset$, then $\{x,y\}\neq\{a^{\prime},b^{\prime}\}$ and $V(\Theta_{f^{\prime}})=\{a^{\prime},b^{\prime},u,v\}$.
Proof.
It is clear that $\{x,y\}\neq\{a^{\prime},b^{\prime}\}$ since $xy\notin E(G)$. Without loss of generality, assume that $y=a^{\prime}$ and let $f^{\prime}$ lie in the pseudo face $F_{2}$. Suppose that $V(\Theta_{f^{\prime}})=\{a^{\prime},b^{\prime},a^{\prime\prime},b^{\prime\prime}\}$. If $V(\Theta_{f^{\prime}})\neq\{a^{\prime},b^{\prime},u,v\}$ (say $a^{\prime\prime}\notin\{u,v\}$), then $f^{\prime}\neq f$. Since $f^{\prime}$ lies in the pseudo face $F_{2}$, it follows that $a^{\prime}b^{\prime}a^{\prime\prime}a^{\prime}\cup\{au,av,uv,bu,bv\}$ is a copy of $C_{3}\dot{\cup}\Theta_{4}$, a contradiction. ∎
Since $x,y$ belong to $\partial(F_{1})$ and $\partial(F_{2})$, respectively, it follows that for each $g\in B_{e}-f$, $\Theta_{e}\cup\Theta_{g}$ is not a copy of $D_{2}$. Further, $\Theta_{e}\cup\Theta_{g}$ cannot be isomorphic to $D_{3}$; otherwise, $\Theta_{e}\cup\Theta_{g}\cup\Theta_{f}$ would contain a copy of $C_{3}\dot{\cup}\Theta_{4}$, a contradiction. Therefore, by combining this with Claim 5.4, we conclude that for each $g\in B_{e}$, $\Theta_{e}\cup\Theta_{g}$ is a copy of $D_{1}$.
Claim 5.5.
The following properties hold.
  1. 1.
    $B_{e}$ is a matching.
  2. 2.
    No edge $g\in B^{*}_{e}$ satisfies: one endpoint of $g$ belongs to $\{u,v\}$ and the other endpoint does not belong to $V(\Theta_{e})$.
Proof.
We prove the first statement. Suppose, to the contrary, that $B_{e}$ is not a matching. Without loss of generality, let $g=bc$ be an edge in $B_{e}$ distinct from $f$ where $f$ and $g$ share the vertex $b$. Since $B_{e}$ contains at least 4 independent edges including $f$, there must be an edge $h\in B_{e}$ such that $h$ have no endpoints in $\{a,b,c\}$ (say $h=pq$). Since $\Theta_{f}\cup\Theta_{e}$, $\Theta_{g}\cup\Theta_{e}$ and $\Theta_{h}\cup\Theta_{e}$ are copies of $D_{1}$, it follows that $\Theta_{xb}\cup ypqy$ is a copy of $C_{3}\dot{\cup}\Theta_{4}$, a contradiction.
Now we show the second statement. Assume, for contradiction, that $g=uc$ is such an edge with $c\notin V(\Theta_{e})$. Since the two faces incident with $g$ are $3$-faces, it follows that $v\notin V(\Theta_{g})$. Since $B_{e}$ contains at least 4 independent edges, there is an edge $h\in B_{e}$ such that $V(\Theta_{g})\cap V(h)=\emptyset$ (say $h=pq$). Hence, $\Theta_{g}\cup vpqv$ is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. ∎
From Claim 5.5, $B_{e}$ is a matching with $B^{*}_{e}\subseteq G[V(\Theta_{e})]$, thus $|B_{e}|\leq(n-2)/2$ and $|B^{*}_{e}|\leq e(G[V(\Theta_{e})])$. Since $xy\notin E(G)$, we have $|B^{*}_{e}|\leq 5$. Consequently, $|E_{I}(G)|\leq|B_{e}|+|B^{*}_{e}|\leq\left\lfloor\frac{n}{2}\right\rfloor+4$, with equality only if $B_{e}$ is a matching of size $\left\lfloor\frac{n-2}{2}\right\rfloor$. Because $\Theta_{e}\cup\Theta_{g}$ forms a $D_{1}$ for each $g\in B_{e}$, $uv+B_{e}$ is a subgraph of $G$. Therefore, if $|E_{I}(G)|=\left\lfloor\frac{n}{2}\right\rfloor+4$, $G$ contains a subgraph isomorphic to $\left(\left\lfloor\frac{n-2}{2}\right\rfloor K_{2}\right)+K_{2}$.
Case 2. For every edge $e$ in $E_{I}(G)$, there is no edge $f$ in $A_{e}$ such that $\Theta_{e}\cup\Theta_{f}$ forms a copy of $D_{1}$.
We next prove that $|E_{I}(G)|<\frac{n}{2}+4$.
Case 2.1. An edge $f\in A_{e}$ exists such that $\Theta_{e}\cup\Theta_{f}$ is a copy of $D_{2}$.
Assume $f=ab$, $e=uv$, $V(\Theta_{e})=\{u,v,x,y\}$, then $V(\Theta_{f})=\{a,b,x,y\}$ as illustrated in Figure 9 (b). We claim $xy\notin E_{I}(G)$. Otherwise, if $xy\in E_{I}(G)$, then since $A_{e}$ contains at least four independent edges, we can choose $h$ such that $h$ and $\Theta_{xy}$ are vertex-disjoint. Thus, $\Theta_{xy}\cup\Theta_{h}$ forms a $D_{1}$, contradicting Case 2. We now show that $A_{e}$ is a matching in $G$. It is enough to prove that any edge $g\in A_{e}$, where $g\neq f$, shares no vertex with $f$. We proceed by contradiction. Assume, without loss of generality, that $g=bq$ (where $b$ is the common vertex of $f$ and $g$). Thus, $g$ is in the byuxbbyuxb pseudo face of $\Theta_{e}\cup\Theta_{f}$. As $|V(\Theta_{e})\cap V(\Theta_{g})|=2$, $\Theta_{e}\cup\Theta_{f}\cup\Theta_{g}$ is isomorphic to $D_{1,1}$ or $D_{1,2}$ (Figure 10). But $D_{1,1}$ and $D_{1,2}$ both contain $C_{3}\dot{\cup}\Theta_{4}$, a contradiction. Hence, $A_{e}$ is a matching.
If there is an edge $e^{\prime}\in A_{e}$ with $e_{1}=a_{1}b_{1}$ such that $\Theta_{e_{1}}\cup\Theta_{e}\cong D_{3}$. As $A_{e}$ is a matching, either $a_{1}b_{1}ya_{1}\cup\Theta_{f}$ or $a_{1}b_{1}xa_{1}\cup\Theta_{f}$ is a copy of $C_{3}\dot{\cup}\Theta_{4}$, a contradiction. Therefore, we have that for each edge $e^{\prime}\in A_{e}$, $\Theta_{e^{\prime}}\cup\Theta_{e}\cong D_{2}$.
Claim 5.6.
$|A_{e}^{*}|\leq 5$.
Proof.
[…]
Next, we complete the proof by considering two separated cases. If $|S|=0$, then $A_{e}^{*}\subseteq E(G[V(\Theta_{e})]-\{xy\})$, implying $|A_{e}^{*}|\leq 5$ (recall that $xy\notin E_{I}(G)$). If $1\leq|S|\leq 2$, then $xy\in E(G)$ and xqyxxqyx is a $3$-face of $G$. We consider the plane graph $D=\Theta_{e}\cup\Theta_{f}\cup\{xy\}$ below. Note that either uxyuuxyu or bxybbxyb is a pseudo face of $D$ (without loss of generality, assume $F=uxyx$ is a pseudo face of $D$). We claim that $ux,uy\notin E_{I}(G)$. Clearly, since $xy\notin E_{I}(G)$ and xqyxxqyx is a $3$-face of $G$, it follows that uxyuuxyu is not a $3$-face of $G$. Therefore, if $ux\in E_{I}(G)$ or $uy\in E_{I}(G)$, then $\Theta_{ux}\cup yaby$ or $\Theta_{uy}\cup xabx$ is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Therefore, $A_{e}^{*}\subseteq S\cup\{uv,vx,vy\}$, implying $|A_{e}^{*}|\leq 5$. ∎
By above discussion, we have that $A_{e}$ forms a matching and $|A_{e}^{*}|\leq 5$. Therefore, $|E_{I}(G)|=|A_{e}^{*}|+|A_{e}|\leq\left\lfloor\frac{n-4}{2}\right\rfloor+5<\left\lfloor\frac{n}{2}\right\rfloor+4$.
eeffgguuvvaabbqqxxyy
(a) $D_{1,1}$
eeffgguuvvqqaabbxxyy
(b) $D_{1,2}$
Figure 10: The plane graphs $D_{1,1}$ and $D_{1,2}$.
Case 2.2. For every edge $e$ in $E_{I}(G)$ and $f\in A_{e}$, $\Theta_{e}\cup\Theta_{f}$ is a copy of $D_{3}$.
Since $A_{e}$ contains at least 4 independent edges, we choose two of them, say $f,g$. It is clear that $\Theta_{e}\cup\Theta_{f}$ and $\Theta_{e}\cup\Theta_{g}$ are copies of $D_{3}$. Hence, $\Theta_{f}\cup\Theta_{g}$ is a copy of $D_{1}$, a contradiction. ∎
Proof of Theorem 1.4: Since
$2e(G)=\sum_{i}if_{i}(G)\geq 3f_{3}(G)+4(f(G)-f_{3}(G))=3f_{3}(G)+4(e(G)+2-n-f_{3}(G)),$
it follows that
$2e(G)\leq f_{3}(G)+4n-8.$ (5)
Let
[…]
Then $E^{\prime}\cap E_{I}(G)=\emptyset$ and thus $|E^{\prime}|\leq e(G)-|E_{I}(G)|$. Since $f_{3}(G)=(|E^{\prime}|+2|E_{I}(G)|)/3$, we have $f_{3}(G)\leq(e(G)+|E_{I}(G)|)/3$. Combining these inequalities, we obtain
$e(G)\leq\left\lfloor\frac{|E_{I}(G)|}{5}+\frac{12n}{5}-\frac{24}{5}\right\rfloor.$ (6)
If $|E_{I}(G)|\leq\frac{n}{2}$, then $e(G)<\left\lfloor\frac{5n}{2}\right\rfloor-4.$ Hence, we assume that $|E_{I}(G)|>\frac{n}{2}$ below.
Claim 5.7.
If $\Delta_{I}(G)\geq 10$ (say $d_{I}(u)=\Delta_{I}(G)$), then GuG-u is $C_{3}$-free. Moreover, $f_{3}(G)\leq n-1$.
Proof.
Let $\Delta_{I}(G)=s$ and $d_{I}(u)=s$. Then $s\geq 10$. Suppose that $d_{G}(u)=t$ and $N_{G}(u)=\{u_{0},u_{1},\ldots,u_{t-1}\}$, where the vertices $u_{0},u_{1},\ldots,u_{t-1}$ are listed in clockwise order around $u$. We further let $E_{I}(u)=\{uu_{c_{0}},uu_{c_{1}},\ldots,uu_{c_{s-1}}\}$. We first prove that every $C=C_{3}$ in $G$ must contain $u$. Suppose, for contradiction, that there exists a 3-cycle $C$ containing no $u$. Then, $|V(C)\cap\{u_{c_{i}}\mid 0\leq i\leq s-1\}|\leq 3$. Since $s\geq 10$, there exists an index $j\in\{0,1,\ldots,s-1\}$ such that $V(C)\cap\{u_{c_{j}-1},u_{c_{j}},u_{c_{j}+1}\}=\emptyset,$ where the subscripts are taken modulo $t$. Consequently, the union $C\cup\Theta_{uc_{j}}$ forms a $C_{3}\dot{\cup}\Theta_{4}$, which is a contradiction. Hence, GuG-u is $C_{3}$-free. Furthermore, every $3$-face of $G$ contains $u$, implying $f_{3}(G)\leq n-1$. ∎
If $\Delta_{I}(G)\geq 10$, then by this Claim 5.7 and Ineq. (5), we have that
$e(G)\leq\frac{f_{3}(G)}{2}+2n-4\leq\frac{5n-1}{2}-4\leq\left\lfloor\frac{5n}{2}\right\rfloor-4.$ (7)
If $\Delta_{I}(G)\leq 9$, then since $n\geq 174$ and $|E_{I}(G)|>\frac{n}{2}$, by lemma 5.3 and Ineq. (6), we obtain
$e(G)\leq\left\lfloor\frac{|E_{I}(G)|}{5}+\frac{12n}{5}-\frac{24}{5}\right\rfloor\leq\left\lfloor\frac{5n}{2}\right\rfloor-4.$ (8)
Therefore, $ex_{\mathcal{P}}(n,C_{3}\dot{\cup}\Theta_{4})\leq\lfloor\frac{5n}{2}\rfloor-4$.
To demonstrate tightness, we now characterize all the extremal graphs. Let $G_{E}$ be an $C_{3}\dot{\cup}\Theta_{4}$-free plane graph with the maximum number of edges. We will describe the characterization of $G_{E}$ under two different circumstances: $\Delta_{I}(G_{E})\leq 9$ and $\Delta_{I}(G_{E})\geq 10$.
If $\Delta_{I}(G_{E})\leq 9$, then by Ineq. (8), $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ if and only if $|E_{I}(G_{E})|=\frac{n}{2}+4$. By Lemma 5.3, we conclude that $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ implies $G_{E}$ contains a spanning subgraph $G^{\prime}$ that is a copy of $\left(\left\lfloor\frac{n-2}{2}K_{2}\right\rfloor\right)+K_{2}$. Without loss of generality, assume that $G^{\prime}=xy+M$, where $M$ is a matching of size $\lfloor(n-2)/2\rfloor$. If $n$ is even, then $e(G^{\prime})=\lfloor\frac{5n}{2}\rfloor-4=e(G)$, and hence $G=G^{\prime}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$. If $n$ is odd, then $|V(G)-V(G^{\prime})|=1$ and $|E(G)-E(G^{\prime})|=2$ (say $V(G)-V(G^{\prime})=\{z\}$ and $E(G)-E(G^{\prime})=\{e_{1},e_{2}\}$). It is clear neither $e_{1}$ nor $e_{2}$ belongs to $G[V(G^{\prime})]$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Hence, we can assume that $e_{1}=z_{1}z$ and $e_{2}=z_{2}z$. Then $\{z_{1},z_{2}\}=\{x,y\}$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Therefore, $G=G^{\prime}\cup\{z_{1}z,z_{2}z\}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$. On the other hand, $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$ is $\Theta_{4}\dot{\cup}C_{3}$-free obviously.
If $\Delta_{I}(G_{E})\geq 10$, then $e(G_{E})<\lfloor\frac{5n}{2}\rfloor-4$ when $n$ is even. Hence, we consider the case where $\Delta_{I}(G_{E})\geq 10$ and $n$ is odd. By the discussion above, $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ if and only if $G_{E}$ satisfies the following two conditions.
  1. 1.
    There exists a vertex $u\in V(G_{E})$ that belongs to n1n-1 3-faces of $G_{E}$, and $f_{3}(G_{E})=n-1$;
  2. 2.
    $n$ is odd and GuG-u is a $C_{3}$-free outerplanar graph with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
Therefore, $G_{E}=u+O$, where $n=|G_{E}|$ is even and $O=G_{E}-u$ is a $C_{3}$-free outerplanar graph with $ex_{\mathcal{OP}}(n,C_{3})=\lfloor\frac{3n}{2}\rfloor-5$ edges. On the other hand, for any planar graph $G=u+O$ with $O=$ a $C_{3}$-free outerplanar graph, we can easily to verify that $G$ is $\Theta_{4}\dot{\cup}C_{3}$-free.
6 Proof of Theorem 1.5
Let $\mathcal{G}$ be the family of maximal $K_{1}+H$-free plane graphs, where $H\in\{P_{2}\dot{\cup}P_{4},2P_{3}\}$. We first present several preliminary lemmas.
Lemma 6.1.
There is a $G\in\mathcal{G}$ such that $G$ is 2-connected.
Proof.
We claim that $G$ is connected; otherwise, adding a bridge would create a $(K_{1}+H)$-free planar graph with more edges, a contradiction. Assume, to the contrary, that every graph in $\mathcal{G}$ contains cut-vertices. We chose an extremal graph $G$ in $\mathcal{G}$ such that the maximum ordered block $D^{*}$ is as large as possible. Since $G$ is not $2$-connected, a vertex $v\in V(D^{*})$ is a cut-vertex of $G$.
It is clear that $D^{*}$ is $2$-connected; otherwise $G$ would be a tree, contradicting the maximality of $e(G)$. Hence, the boundary of $D^{*}$ is a cycle, denoted by $C^{*}$. Let $u$ be a neighbor of $v$ in $C^{*}$. It is clear that the block-tree of $D^{*}-\{uv\}$ a path such that
  1. 1.
    $u,v$ belong to the two leaf blocks, respectively.
  2. 2.
    neither $u$ nor $v$ is a cut-vertex of $D^{*}-\{uv\}$.
Because $v$ is a cut-vertex of $G$, there exists a block $D\neq D^{*}$ in $G$. Choose a vertex $w\neq v$ on the boundary of $D$, and let $P$ be a $vw$-path in $D$. Construct $G^{\prime}=G-\{uv\}+\{uw\}$. The edge $uv$ is not in any triangle in $G^{\prime}$, and since every edge in $K_{1}+H$ is in a triangle, $G^{\prime}$ remains $K_{1}+H$-free. Since $e(G)=e(G^{\prime})$, $G^{\prime}$ is also an extremal plane graph in $\mathcal{G}$. However, $R=(D^{*}-\{uv\})\cup P\cup\{uw\}$ is 2-connected in $G^{\prime}$ and $|R|>|D^{*}|$. Thus, $G^{\prime}$ contains a block strictly larger than $D^{*}$, contradicting the maximality of $D^{*}$. Consequently, $G$ is $2$-connected. ∎
Lemma 6.2.
Let $G$ be a 2-connected plane graph that is either $(K_{1}+(P_{2}\cup P_{4}))$-free or $(K_{1}+2P_{3})$-free. Then for any vertex $v$ with $d(v)\geq 6$ in $G$, $v$ is adjacent to at least $\frac{d(v)}{3}-1/3$ non-3-faces.
Proof.
Let $v$ be a vertex with $d(v)\geq 6$. We consider two cases based on the forbidden subgraph condition:
Case 1: $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free.
Since $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free, each induced path in $G[N(v)]$ has order at most 5. If $G[N(v)]$ contains a path $P$ with $4\leq|P|\leq 5$, then $G[N(v)\setminus V(P)]$ must be a star. This ensures that $v$ is incident to at least $d(v)-|P|+1\geq\frac{d(v)}{3}$ non-3-faces for $d(v)\geq 6$. If every induced path in $G[N(v)]$ has order at most 3, then $v$ is trivially incident to at least $\frac{d(v)}{3}$ non-3-faces.
Case 2: $G$ is $(K_{1}+2P_{3})$-free.
For this case, each path in $G[N(v)]$ has order at most 5. If $G[N(v)]$ contains a path $P$ with $3\leq|P|\leq 5$, then the remaining graph $G[N(v)\setminus E(P)]$ can only contain paths of order at most 2. This structure guarantees that $v$ is incident to at least $\lceil\frac{d(v)-|P|}{2}\rceil+1\geq\frac{d(v)}{3}-1/3$ non-3-faces for $d(v)\geq 6$. If all induced paths in $G[N(v)]$ have order at most 2, then $v$ is incident to at least $\lceil\frac{d(v)}{2}\rceil\geq\frac{d(v)}{3}$ non-3-faces. ∎
Proof of Theorem 1.5 For the lower bound, we construct a $(K_{1}+H)$-free plane graph on $n$ vertices with $\frac{11n}{4}-3$ edges. Let $n=12k$ for an integer $k\geq 1$. We define the base graph $Q_{1}$ as shown in Figure 11(a). For $k\geq 2$, the graph $Q_{k}$ is constructed recursively by embedding $Q_{k-1}$ into the central triangle of the structure in Figure 11(b), such that the three vertices on the center triangle and the three vertices on the outer boundary of $Q_{k-1}$ are connected by three edges that forms a matching. It is straightforward to verify that each $Q_{k}$ is $(K_{1}+H)$-free and contains exactly $\frac{11n}{4}-3$ edges.
(a) $Q_{1}$
$F_{k-1}$
(b) $Q_{k}$
Figure 11: The construction of $F_{k}$.
We now establish the upper bound. By Lemma 6.1, we may assume $G\in\mathcal{G}$ is 2-connected. Applying a discharging argument to bound the number of missing chords in $D$, we initially assign each $k$-face a charge of k3k-3. This charge is then equally distributed to its incident vertices, yielding at least $\frac{k-3}{k}\geq\frac{1}{4}$ per vertex. Lemma 6.2 further guarantees that each vertex $v$ with $d(v)\geq 6$ is incident to at least $\frac{d(v)}{3}-1/3$ non-3-faces, enabling it to receive a total charge of at least $\frac{1}{4}(\frac{d(v)}{3}-1/3)$. Consequently, we derive
$3n-6-e(G)\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left\lfloor\frac{d(v)}{3}\right\rfloor\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left(\frac{d(v)}{3}-1/3\right).$
Furthermore, since $5x+\sum_{\{v:d(v)\geq 6\}}d(v)\geq 2e(G)$, where $x$ denotes the number of vertices of degree at most 5, we obtain:
$2e-5x\leq\sum_{\{v:d(v)\geq 6\}}d(v)\leq 36n-72-12e(G)+(n-x).$
Hence, $e(G)\leq\frac{37n-72+4x}{14}\leq\frac{41n}{14}-\frac{36}{7}$, as desired. $\Box$
Based on the construction for the lower bound in Theorem 1.5, a natural question arises:
Question. Is the lower bound in Theorem 1.5 tight when $14|n$?
7 Acknowledgments
Ping Li is suppose by National Natural Science Foundation of China (No. 12201375); Xuqing Bai is supported by the Fundamental Research Funds for the Central Universities (No. ZYTS24069); Zhipeng Gao is supported by the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-001) and the Fundamental Research Funds for the Central Universities (No. ZYTS24076).

### v2

@@
 One representative graph among such union graphs is $K_{1}+L$ where $L$ is a linear forest without isolated vertices.
-Previous works solved the cases when $L$ is a path or a matching.
+Previous works solved the cases in which $L$ is a path or a matching, or satisfies $|L|\geq 7$.
 In this work, we first investigate the planar Turán number of the graph $K_{1}+L$ when $L$ is the disjoint union of a $P_{2}$ and $P_{3}$.
@@
 Furthermore, we completely characterize all extremal graphs for the remaining two of these three cases.
-Additionally, for two other unsolved cases involving $K_{1}+L$, we establish improved bounds.
 Keywords: Planar Turán number, small graphs, union of triangles, extremal graphs
@@
 Given the aforementioned challenges, researchers have shifted their focus toward determining the exact planar Turán numbers of small planar graphs.
-So far, tight bounds are known only for a few small planar graphs, including short cycles of length up to eight [1, 5, 6], paths of order at most eleven [8], and small $\Theta$-graphs [10] (we use $\Theta_{k}$ to denote a set of graphs that are obtained from a cycle by adding an additional edge joining two non-consecutive vertices.
+So far, tight bounds are known only for a few small planar graphs, including short cycles of length up to seven [1, 5, 6], paths of order at most eleven [8], and small $\Theta$-graphs [10] (we use $\Theta_{k}$ to denote a set of graphs that are obtained from a cycle by adding an additional edge joining two non-consecutive vertices.
 It is clear that $\Theta_{k}$ is a single graph if $k=4,5$).
@@
 Among these configurations, one of the most interesting is $K_{1}+L_{t}$, where $L_{t}$ is a linear forest of $t$ vertices (the graph $K_{1}+L_{t}$ is obtained by joining a new vertex to all the vertices of $L_{t}$).
-Lan, Shi, and Song [9] proved that
-Theorem 1.1 ([9]).
-For an integer $4\leq t\leq 6$, let $H$ be a graph on $t$ vertices consisting of disjoint paths. Then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{13(t-1)n}{4t-2}-\frac{12(t-1)}{2t-1}$ for all $n\geq t+1$.
+Lan, Shi, and Song [9] proved that ffor an integer $4\leq t\leq 6$, let $H$ be a graph on $t$ vertices consisting of disjoint paths.
+Then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{13(t-1)n}{4t-2}-\frac{12(t-1)}{2t-1}$ for all $n\geq t+1$.
 It is worth noting that not all the upper bounds obtained are tight.
@@
 Because $H_{1}$ is a subgraph of $H_{3}$, $ex_{\mathcal{P}}(n,H_{3})=3n-6$ follows directly.
-In addition to $H_{4}$, we also focus on determining the planar Turán numbers of the remaining two graphs $H_{5}$, and $H_{6}$.
+In addition to $H_{4}$, we also focus on determining the planar Turán numbers of the remaining two graphs $H_{5}$ and $H_{6}$.
 For $H_{5}$, we determine its planar Turán number exactly when $n=10x+6y$ has integer solutions $x\geq 2$ and $y\geq 0$, and characterize all extremal graphs.
@@
 We determine the planar Turán number of $H_{6}$, which is the disjoint union of $C_{3}$ and $\Theta_{4}$, and characterize all extremal graphs.
-For convenience, we use $M_{t}$ to denote the union of $\left\lfloor\frac{t}{2}\right\rfloor$ pairwise vertex-disjoint edges and $\left\lceil\frac{t}{2}\right\rceil-\left\lfloor\frac{t}{2}\right\rfloor$ isolated vertices.
-To describe the extremal $H_{6}$-free planar graphs, we need the outerplanar Turán number of $C_{3}$ (denoted $ex_{\mathcal{OP}}(n,C_{3})$), which is the maximum number of edges in an $n$-vertex $C_{3}$-free outerplanar graph.
+For this purpose, we first introduce two graphs.
+We use $M_{t}$ to denote the union of $\left\lfloor\frac{t}{2}\right\rfloor$ pairwise vertex-disjoint edges and $\left\lceil\frac{t}{2}\right\rceil-\left\lfloor\frac{t}{2}\right\rfloor$ isolated vertices.
+For odd $n$, let $K_{2}\vee M_{n-2}$ denote the graph obtained from $K_{2}+M_{n-3}$ by adding an additional vertex $u$ and two edges $uv_{1},uv_{2}$, where $v_{1}$ and $v_{2}$ are endpoints of two arbitrary edges in $M_{n-2}$, respectively.
+To describe the extremal $H_{6}$-free planar graphs, we also need the notion outerplanar Turán number of $C_{3}$ (denoted $ex_{\mathcal{OP}}(n,C_{3})$), which is the maximum number of edges in an $n$-vertex $C_{3}$-free outerplanar graph.
 Fang and Zhai [4] proved that $ex_{\mathcal{OP}}(n,C_{3})=\frac{3n-4}{2}$ for each even integer $n\geq 4$.
@@
 If $n\geq 174$, then $ex_{\mathcal{P}}(n,C_{3}\dot{\cup}\Theta_{4})=\left\lfloor\frac{5n}{2}\right\rfloor-4$.
-Moreover, if $n$ is even, then the extremal planar graph is a copy of $M_{n-2}+K_{2}$; if $n$ is odd, then the extremal planar graph is either a copy of $M_{n-2}+K_{2}$ or $O+\{u\}$, where $O$ represents a $C_{3}$-free outerplanar graph of even order with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
-Let $L^{\prime}_{k}$ denote a linear forest of order $k$ without isolated vertices.
-Authors in [1, 3, 9] determined $ex_{\mathcal{P}}(n,K_{1}+L^{\prime}_{k})$ for all $L^{\prime}_{k}$ except for the two cases $L^{\prime}_{6}=P_{2}\dot{\cup}P_{4}$ and $L^{\prime}_{6}=2P_{3}$.
-We give a lower bound and an upper bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H\in\{P_{2}\cup P_{4},2P_{3}\}$, which offers a slight improvement to the upper bounds provided by Lan, Shi and Song (Theorem 1.1) for these scenarios.
-Theorem 1.5.
-Let $H\in\{P_{2}\cup P_{4},2P_{3}\}$, then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{41n}{14}-\frac{36}{7}$. In addition, if $12|n$, then $ex_{\mathcal{P}}(n,K_{1}+H)\geq\frac{11n}{4}-3$.
+Moreover, if $n$ is even, then the extremal planar graph is a copy of $M_{n-2}+K_{2}$; if $n$ is odd, then the extremal planar graph is either a copy of $K_{2}+M_{n-2}$, or $K_{2}\vee M_{n-2}$, or $\{u\}+O$, where $O$ represents a $C_{3}$-free outerplanar graph of even order with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
 The rest of this paper is organized as follows.
@@
 In Sections 3, 4 and 5, we establish sharp bounds on the planar Turán number of $H_{4},H_{5}$ and $H_{6}$, respectively, and characterize all extremal planar graphs for the latter two graphs.
-In Section 6, we give a short proof of Theorem 1.5.
 2 Preliminaries
@@
 We call such a new TB a solid TB (i.e., a solid TB contains no holes with facial cycle $C_{3}$).
-In the proofs of Theorems 1.2 and 1.3, we will estimate the maximum triangle-density among all $H$-free TBs and TCs (where $H$ is either $H_{4}$ or $H_{5}$).
+In the proofs of Theorems 1.1 and 1.2, we will estimate the maximum triangle-density among all $H$-free TBs and TCs (where $H$ is either $H_{4}$ or $H_{5}$).
 For this purpose, it suffices to estimate the maximum triangle-density among all $H$-free solid TBs and all $H$-free TCs composed of solid TBs.
@@
 Hence, we can assume that $e_{1}=z_{1}z$ and $e_{2}=z_{2}z$.
-Then $\{z_{1},z_{2}\}=\{x,y\}$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction.
-Therefore, $G=G^{\prime}\cup\{z_{1}z,z_{2}z\}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$.
-On the other hand, $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$ is $\Theta_{4}\dot{\cup}C_{3}$-free obviously.
+Then either $\{z_{1},z_{2}\}=\{x,y\}$ or $z_{1},z_{2}$ are endpoints of two edges in the matching $G^{\prime}-\{x,y\}$, respectively, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction.
+Therefore, $G=G^{\prime}\cup\{z_{1}z,z_{2}z\}$ is either a copy of $K_{2}+\left(\frac{n-2}{2}K_{2}\right)=K_{2}+M_{n-2}$ or a copy of $K_{2}\vee M_{n-2}$.
+On the other hand, $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$ and $K_{2}\vee M_{n-2}$ are $\Theta_{4}\dot{\cup}C_{3}$-free obviously.
 If $\Delta_{I}(G_{E})\geq 10$, then $e(G_{E})<\lfloor\frac{5n}{2}\rfloor-4$ when $n$ is even.
@@
 On the other hand, for any planar graph $G=u+O$ with $O=$ a $C_{3}$-free outerplanar graph, we can easily to verify that $G$ is $\Theta_{4}\dot{\cup}C_{3}$-free.
-6 Proof of Theorem 1.5
-Let $\mathcal{G}$ be the family of maximal $K_{1}+H$-free plane graphs, where $H\in\{P_{2}\dot{\cup}P_{4},2P_{3}\}$. We first present several preliminary lemmas.
-Lemma 6.1.
-There is a $G\in\mathcal{G}$ such that $G$ is 2-connected.
-Proof.
-We claim that $G$ is connected; otherwise, adding a bridge would create a $(K_{1}+H)$-free planar graph with more edges, a contradiction.
-Assume, to the contrary, that every graph in $\mathcal{G}$ contains cut-vertices.
-We chose an extremal graph $G$ in $\mathcal{G}$ such that the maximum ordered block $D^{*}$ is as large as possible.
-Since $G$ is not $2$-connected, a vertex $v\in V(D^{*})$ is a cut-vertex of $G$.
-It is clear that $D^{*}$ is $2$-connected; otherwise $G$ would be a tree, contradicting the maximality of $e(G)$.
-Hence, the boundary of $D^{*}$ is a cycle, denoted by $C^{*}$.
-Let $u$ be a neighbor of $v$ in $C^{*}$.
-It is clear that the block-tree of $D^{*}-\{uv\}$ a path such that
-  1. 1.
-    $u,v$ belong to the two leaf blocks, respectively.
-  2. 2.
-    neither $u$ nor $v$ is a cut-vertex of $D^{*}-\{uv\}$.
-Because $v$ is a cut-vertex of $G$, there exists a block $D\neq D^{*}$ in $G$.
-Choose a vertex $w\neq v$ on the boundary of $D$, and let $P$ be a $vw$-path in $D$.
-Construct $G^{\prime}=G-\{uv\}+\{uw\}$.
-The edge $uv$ is not in any triangle in $G^{\prime}$, and since every edge in $K_{1}+H$ is in a triangle, $G^{\prime}$ remains $K_{1}+H$-free.
-Since $e(G)=e(G^{\prime})$, $G^{\prime}$ is also an extremal plane graph in $\mathcal{G}$.
-However, $R=(D^{*}-\{uv\})\cup P\cup\{uw\}$ is 2-connected in $G^{\prime}$ and $|R|>|D^{*}|$.
-Thus, $G^{\prime}$ contains a block strictly larger than $D^{*}$, contradicting the maximality of $D^{*}$.
-Consequently, $G$ is $2$-connected.
-∎
-Lemma 6.2.
-Let $G$ be a 2-connected plane graph that is either $(K_{1}+(P_{2}\cup P_{4}))$-free or $(K_{1}+2P_{3})$-free.
-Then for any vertex $v$ with $d(v)\geq 6$ in $G$, $v$ is adjacent to at least $\frac{d(v)}{3}-1/3$ non-3-faces.
-Proof.
-Let $v$ be a vertex with $d(v)\geq 6$. We consider two cases based on the forbidden subgraph condition:
-Case 1: $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free.
-Since $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free, each induced path in $G[N(v)]$ has order at most 5.
-If $G[N(v)]$ contains a path $P$ with $4\leq|P|\leq 5$, then $G[N(v)\setminus V(P)]$ must be a star.
-This ensures that $v$ is incident to at least $d(v)-|P|+1\geq\frac{d(v)}{3}$ non-3-faces for $d(v)\geq 6$.
-If every induced path in $G[N(v)]$ has order at most 3, then $v$ is trivially incident to at least $\frac{d(v)}{3}$ non-3-faces.
-Case 2: $G$ is $(K_{1}+2P_{3})$-free.
-For this case, each path in $G[N(v)]$ has order at most 5.
-If $G[N(v)]$ contains a path $P$ with $3\leq|P|\leq 5$, then the remaining graph $G[N(v)\setminus E(P)]$ can only contain paths of order at most 2.
-This structure guarantees that $v$ is incident to at least $\lceil\frac{d(v)-|P|}{2}\rceil+1\geq\frac{d(v)}{3}-1/3$ non-3-faces for $d(v)\geq 6$.
-If all induced paths in $G[N(v)]$ have order at most 2, then $v$ is incident to at least $\lceil\frac{d(v)}{2}\rceil\geq\frac{d(v)}{3}$ non-3-faces.
-∎
-Proof of Theorem 1.5 For the lower bound, we construct a $(K_{1}+H)$-free plane graph on $n$ vertices with $\frac{11n}{4}-3$ edges.
-Let $n=12k$ for an integer $k\geq 1$.
-We define the base graph $Q_{1}$ as shown in Figure 11(a).
-For $k\geq 2$, the graph $Q_{k}$ is constructed recursively by embedding $Q_{k-1}$ into the central triangle of the structure in Figure 11(b), such that the three vertices on the center triangle and the three vertices on the outer boundary of $Q_{k-1}$ are connected by three edges that forms a matching.
-It is straightforward to verify that each $Q_{k}$ is $(K_{1}+H)$-free and contains exactly $\frac{11n}{4}-3$ edges.
-(a) $Q_{1}$
-$F_{k-1}$
-(b) $Q_{k}$
-Figure 11: The construction of $F_{k}$.
-We now establish the upper bound.
-By Lemma 6.1, we may assume $G\in\mathcal{G}$ is 2-connected.
-Applying a discharging argument to bound the number of missing chords in $D$, we initially assign each $k$-face a charge of k3k-3.
-This charge is then equally distributed to its incident vertices, yielding at least $\frac{k-3}{k}\geq\frac{1}{4}$ per vertex.
-Lemma 6.2 further guarantees that each vertex $v$ with $d(v)\geq 6$ is incident to at least $\frac{d(v)}{3}-1/3$ non-3-faces, enabling it to receive a total charge of at least $\frac{1}{4}(\frac{d(v)}{3}-1/3)$.
-Consequently, we derive
-$3n-6-e(G)\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left\lfloor\frac{d(v)}{3}\right\rfloor\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left(\frac{d(v)}{3}-1/3\right).$
-Furthermore, since $5x+\sum_{\{v:d(v)\geq 6\}}d(v)\geq 2e(G)$, where $x$ denotes the number of vertices of degree at most 5, we obtain:
-$2e-5x\leq\sum_{\{v:d(v)\geq 6\}}d(v)\leq 36n-72-12e(G)+(n-x).$
-Hence, $e(G)\leq\frac{37n-72+4x}{14}\leq\frac{41n}{14}-\frac{36}{7}$, as desired. $\Box$
-Based on the construction for the lower bound in Theorem 1.5, a natural question arises:
-Question. Is the lower bound in Theorem 1.5 tight when $14|n$?
 6 Acknowledgments